### v1

# Leveraging Large Language Model for Information Retrieval-based Bug Localization

Abstract
Information Retrieval-based Bug Localization aims to identify buggy source files for a given bug report. While existing approaches—ranging from vector space models to deep learning models—have shown potential in this domain, their effectiveness is often limited by the vocabulary mismatch between bug reports and source code. To address this issue, we propose a novel Large Language Model (LLM) based bug localization approach, called GenLoc. Given a bug report, GenLoc leverages an LLM equipped with code-exploration functions to iteratively analyze the code base and identify potential buggy files. To gather better context, GenLoc may optionally retrieve semantically relevant files using vector embeddings. GenLoc has been evaluated on over 9,000 real-world bug reports from six large-scale Java projects. Experimental results show that GenLoc outperforms five state-of-the-art bug localization techniques across multiple metrics, achieving an average improvement of more than 60% in Accuracy@1.
Index Terms:
bug localization, debugging, large language model, information-retrieval
I Introduction
Bug localization is the process of identifying the faulty or buggy locations of the source code [1]. It is a crucial step in software debugging to ensure the quality of a software [2]. However, large systems often incur a huge number of bugs everyday. For example, everyday almost 300 bugs are reported in Mozilla, which is overwhelming for the developers [3]. Additionally, given the size and complexity of large-scale software systems today, manually locating bugs is time consuming, tedious and expensive task [4]. As a result, there is a strong demand for automated bug localization techniques to ease the burden of developers and speed up the program debugging process [5].
Over the years, researchers have proposed multiple approaches for bug localization. Among these, Spectrum-based Bug Localization (SBBL) and Information Retrieval-based Bug Localization (IRBL) are among the most studied and widely recognized techniques [6, 5]. SBBL leverages runtime behavior—such as test coverage information—to analyze which parts of a program are exercised by passing and failing test cases, and ranks program elements based on their likelihood of being faulty [7]. In contrast, IRBL takes a bug report—a document that describes a software bug—as input and does not require bug-revealing test cases or execution traces [8]. Due to its simplicity, IR-based approaches have received growing attention in recent years.
IRBL approaches consider the bug report as a query and the software repository as a collection of documents. Given a bug report, these approaches output a ranked list of documents (source files) based on their probability of being faulty. The goal is to rank the actual buggy files as high as possible. For calculating the similarity between bug reports and the source files, earlier techniques use Vector Space Model (VSM) [9, 10], Topic Modeling [11, 12], while more recent methods [13, 14, 15, 16, 17, 18, 5, 19, 20, 21] rely on machine learning and deep learning models. Many of these approaches also incorporate additional features such as historical bug-fix data. However, such information may not always be available, especially in new projects, limiting the generalizability of these methods. Furthermore, the effectiveness of these approaches is often limited by the vocabulary gap between bug reports and source code, as they tend to use different terminologies [6].
Recently, Large Language Models (LLMs) have demonstrated promising results in various software engineering tasks including code generation, code summarization and program repair [22, 23]. Their ability to understand natural language and source code might help improve information retrieval-based bug localization. However, applying LLMs to this task presents unique challenges. Most notably, due to context window limitations, it is infeasible to provide the entire code base as input and ask the LLM to identify relevant files for a given bug report. A straightforward alternative—feeding one file at a time to LLM and asking whether it is relevant—quickly becomes impractical due to the large number of files in real-world repositories. Furthermore, LLMs often struggle to identify and leverage relevant information when the input context is long and relevant information is buried in the middle [24]. Another critical challenge is that LLMs are prone to hallucination (e.g., generating non-existent file names) particularly when they lack sufficient contextual information. Therefore, an effective strategy is needed to help LLMs navigate the code base, selectively retrieve relevant information, and reason within a constrained context window.
In this context, we propose a novel bug localization technique, named GenLoc, that combines semantic retrieval with the step-by-step reasoning capabilities of LLMs to identify buggy files for a given bug report. The name GenLoc stands for Generative AI-based Bug Localization, emphasizing its core design principle of leveraging generative language models for fault localization. Unlike many prior approaches [9, 5], GenLoc relies solely on the bug report and does not require any additional project-specific metadata or historical bug-fix data. At first, GenLoc converts the source code into chunks and measures their semantic similarity with the bug report to retrieve a set of potentially relevant files. After analyzing a bug report, an LLM may choose to inspect these retrieved files further or explore other files if needed. To support this exploration, the LLM is equipped with a set of functions (e.g., retrieving method signatures or method bodies) that enable it to iteratively reason about which files are likely related to the reported bug. This approach allows the LLM to conduct more guided and selective localization.
Figure 1: Workflow of GenLoc
GenLoc is evaluated on over 9,000 real-world bug reports from six large open-source Java projects and compared against five state-of-the-art traditional and deep learning-based IRBL techniques. The results show that GenLoc outperforms all existing approaches across multiple metrics, including Accuracy@1 and Accuracy@5. On average, GenLoc achieves an Accuracy@1 of 43.19%, surpassing all prior techniques by at least 60%. Furthermore, GenLoc offers complementary strengths to existing IRBL techniques, suggesting potential for future integration to further improve bug localization performance.
In summary, this paper makes the following contributions:
  • The first study to investigate the impact of incorporating LLMs for information retrieval-based bug localization.
  • The introduction of a novel bug localization approach named GenLoc that integrates semantic retrieval with LLM-guided reasoning.
  • A comprehensive empirical evaluation of GenLoc on over 9,000 real-world bugs from six large open-source Java projects.
  • A publicly accessible replication package [^1] https://github.com/GenLoc-2025/GenLoc to facilitate future research.
II Background
This section outlines the key technical components that form the foundation of the proposed approach.
Embedding: It represents complex data, such as text, images, or source code, as numerical vectors in a high-dimensional space. The core idea is that semantically similar items (e.g., two similar sentences or code snippets) will have embeddings that are close to each other in that space. In GenLoc, both source files and bug reports are converted into embeddings to enable semantic similarity calculation between them.
Vector database: A vector database is a special type of database designed to store and manage large collections of high-dimensional embeddings and perform efficient similarity searches [25]. To accelerate the retrieval process, vector databases commonly use Approximate Nearest Neighbor (ANN) algorithms [26], which find vectors that are close to a given query without exhaustively comparing all entries. In GenLoc, the embeddings of all the source files are stored in a vector database to facilitate retrieval of the top source files that are most semantically similar to a given bug report.
Function calling: This mechanism allows LLMs to interact with external tools and systems to retrieve up-to-date information or perform specific tasks that they cannot directly do. For example, if LLM is given the prompt "What is the current temperature in Los Angeles?", it will fail to answer because it lacks access to real-time data. However, it can be provided with access to a function like get_current_temperature(), which queries an external weather API and returns the latest temperature. The model can then call this function, receive the temperature, and integrate it into its response. In GenLoc, the LLM is given access to several functions such as retrieving method signatures and bodies that allow it to iteratively navigate the code base and identify potentially buggy files.
ReAct framework: The ReAct (Reasoning and Acting) framework enables language models to interleave natural language reasoning with tool-based actions, thereby enhancing their problem-solving capabilities [27]. Instead of generating a final answer in a single step, the model alternates between articulating intermediate reasoning steps and performing actions, such as calling a function. This synergy between reasoning and acting allows the model to iteratively refine its understanding, validate hypotheses, and make informed decisions. Since ReAct has shown promise in handling complex tasks that require multi-step thinking, access to external environments, or dynamic decision-making [27, 28], this paper adopts the ReAct framework to guide the LLM in navigating the code base and progressively narrowing down the search space during bug localization.
III Methodology
GenLoc operates in two primary steps to localize relevant files based on a given bug report. First, it retrieves a set of semantically similar files using embedding-based similarity. Next, an LLM, supported by a set of external functions, iteratively analyzes the bug report and source files. It may inspect the embedding-based retrieved files or explore other parts of the code base to generate a ranked list of potentially buggy files. An overview of the workflow is shown in Fig. 1.
III-A Shortlisting suspicious files
To identify source files potentially related to a given bug report, this step computes the semantic similarity between the bug report and the source code using embedding-based representations. The process begins by traversing all source files in the project. For each file, its fully qualified file path and method bodies are extracted. Each file is then represented as a concatenation of its fully qualified file path and the method bodies. The inclusion of the file path provides contextual information about the file’s role or feature within the overall system architecture (e.g., org.eclipse.jdt.ui/ui/org/eclipse/jdt/ui/JavaElementLabels.java suggests a UI-related component), while method bodies capture the actual implementation logic, which is crucial since methods serve as the fundamental units of behavior in a program.
The concatenation of a file’s path and method bodies can result in lengthy text. To better localize relevant content, this representation is split into smaller chunks if it exceeds a predefined length threshold [29]. Each chunk is then converted into an embedding vector and stored in a vector database. Similarly, the textual content of the bug report—comprising its summary and description—is transformed into embedding vector.
To measure the similarity between the bug report and the source code, cosine similarity is employed, which measures the angular distance between the embedding vectors [30]. Based on the similarity scores, a predefined number of the top suspicious files are shortlisted. These files serve as candidate inputs for further analysis by the LLM in the subsequent step.
III-B Ranking suspicious files
In this step, an LLM is employed to generate a ranked list of potentially buggy source files based on a given bug report, as illustrated in Fig. 1. To enhance the LLM’s reasoning capabilities, the prompt (shown in Fig. 2) is constructed by following the ReAct (Reasoning and Acting) framework [27], which enables LLMs to interleave natural language reasoning with tool-based actions. This approach has proven effective in improving performance on complex decision-making tasks by enabling iterative reasoning and environment interaction.
At first, the LLM is prompted to assume the role of an expert software engineer specializing in fault localization, as role-playing improves the reasoning abilities of LLMs [31]. Next, the prompt is designed to decompose bug localization into a series of smaller tasks since LLMs perform better when complex tasks are broken down into sub-tasks [32]. These sub-tasks include analyzing the bug report, identifying potential file or method names, analyzing method signatures and bodies, and integrating evidence to prioritize suspicious files.
You are an expert software engineer specializing in fault localization. Your goal is to identify the most probable buggy Java files based on a given bug report. You have access to five functions that will help you infer file names, locate methods, and analyze source code. You must follow an iterative, reasoning-based approach, refining your strategy dynamically based on prior successes and failures. Continue this process until you either (a) produce a well-justified ranked list of the 10 most relevant files based on the bug report, or (b) reach the maximum limit of 10 iterations. In the 10th iteration, you must provide your final output regardless of confidence level. Workflow 1. Analyze the Bug Report: Extract relevant keywords, error messages, and functional hints from the bug summary and description. Identify potential components (e.g., UI, database, networking) involved in the issue. 2. Search: Use search_file() to check if a filename matching the extracted keywords or functionality exists in the code base. If the bug report references a specific method name, use search_method() to locate the file(s) containing that method. If an inferred filename or method location does not exist, refine your strategy: adjust assumptions, explore variations, and retry. If no strong inference can be made, use get_candidate_filenames() to retrieve 50 potential filenames. From the retrieved filenames, prioritize those that align with the bug report’s keywords, functionality, or mentioned methods. 3. Method Analysis: For shortlisted files, retrieve method signatures using get_method_signatures_of_a_file(). Identify methods that directly align with the bug’s context (e.g., matching function names, handling related data). If method signatures suggest a relevant function, retrieve its implementation using get_method_body(). Analyze logic to determine if it aligns with the bug’s symptoms. 4. Refinement and Ranking: Rank files based on multiple factors: Keyword and functionality match Method or filename alignment with bug context Code logic alignment with the bug description If uncertainty remains, refine the analysis by iterating over previous steps with adjusted assumptions. 5. Output: Provide a ranked list of the 10 most relevant files based on their likelihood of containing the bug. Ensure filenames exactly match those provided—do not modify case, structure, or abbreviate them. Justify each file’s inclusion, clearly explaining its relevance to the bug.
Figure 2: LLM Prompt
To support the reasoning process, the LLM is given access to the following five external functions. The LLM may invoke one or more of these functions over a fixed number of iterations to get information about the code base.
  1. 1.
    search_file(): This function enables the LLM to check whether a specific file exists in the code base. It can be used to (i) verify the existence of a file explicitly mentioned in the bug report (e.g., checking if JavaElementLabels.java exists when “JavaElementLabels” is referenced), or (ii) infer and validate a possible filename derived from textual clues in the bug report.
  2. 2.
    search_method(): When a method name is mentioned in the bug report, e.g., updateLabel(), this function enables the LLM to find which files contain its definition and thereby narrowing down the search space.
  3. 3.
    get_candidate_filenames(): This function retrieves a list of potentially relevant files based on their semantic similarity to the bug report from the previous step (Section III-A). The LLM can leverage this list as an initial pool of candidates, especially when the bug report does not provide any clues.
  4. 4.
    get_method_signatures_of_a_file(): Once a file is identified, this function enables the model to inspect the file’s high-level structure by retrieving all method signatures defined within it. This helps the LLM determine whether the file’s responsibilities align with the bug report.
  5. 5.
    get_method_body(): To evaluate a method’s implementation, the model can call this function to retrieve the body of a given method. This allows a deeper inspection of whether the method may be related to the described issue.
At each iteration, the model can iteratively “reason” about what information is currently available, determine what additional data is required, and then “act” by invoking one of these functions. The outcome of the function call is integrated into the LLM’s internal reasoning in the next iteration. This iterative loop continues until the model either reaches the maximum allowed number of steps or gathers sufficient information to generate a ranked list of suspicious files.
As LLMs are susceptible to hallucinations, such as generating filenames or method signatures that do not exist in the code base, the system is equipped with recovery strategies to mitigate these issues. For example, when a generated method signature does not match any entry in the code base, approximate candidates are retrieved using fuzzy matching based on the Damerau-Levenshtein distance [33]. Lastly, the LLM is prompted to output a ranked list where each entry includes a fully qualified file path along with a brief justification for its selection. Since multiple files from different packages can have the same name, the fully qualified file path is used to uniquely identify each file. The inclusion of justifications is motivated by prior work demonstrating that self-explanatory prompts significantly improve LLM comprehension [34].
Although several strategies are employed to mitigate LLM hallucinations, such as grounding it through function calls, the model can still generate filenames that do not exist in the code base. To address this, a post-processing step is applied to verify the existence of each predicted file. If the predicted fully qualified file path matches an existing file, it is retained. Otherwise, the base filename (e.g., TitleBlock.java) is extracted, and all files with the same name are retrieved. Among these candidates, the file whose fully qualified file path has the highest Jaccard similarity [35] with the predicted filename is selected. If no suitable match is found, the file is excluded from the final ranked list.
IV Experiement Setup
This section presents the implementation details, research questions, baseline techniques, benchmark dataset, and evaluation metrics.
IV-A Implementation
Similar to most of the existing bug localization techniques [20, 5, 21], GenLoc is implemented in Java since it is one of the most popular programming languages. However, the approach is easily extendable to other languages, as it leverages the Tree-sitter library [36] for source code parsing, which supports more than ten programming languages. To compute semantic similarity between bug reports and source code files, GenLoc employs OpenAI’s text-embedding-3-small model due to its cost-effectiveness (\$0.02 per 1M tokens) [37]. Following established methods [38, 39], the source code is divided into chunks of 300 tokens. For each bug report, the top 50 source files are retrieved based on their embedding similarity to the report.
To improve efficiency and minimize embedding costs, embeddings for all source files are generated and stored in ChromaDB [40] when processing the first bug in each project. Since different bugs occur in different versions, for each subsequent bug, only the embeddings of files that have been added, modified, deleted, or renamed across versions are updated. For inference, GPT-4o mini by OpenAI is used due to its cost-effectiveness (\$0.15 per 1M input tokens and \$0.60 per 1M output tokens) [41]. All model parameters are kept at their default settings to avoid the computational cost of hyperparameter tuning. In particular, the temperature parameter is left at its default value (1.0) (rather than being set to 0) to enable diverse reasoning paths during inference [42, 43]. A temperature of 0 significantly reduces variability by consistently selecting the highest-probability tokens, which can lead the model to overlook less obvious but relevant files. Our preliminary analysis also showed that the performance remained mostly unchanged regardless of the temperature setting. This aligns with prior findings that changes in temperature between 0.0 and 1.0 do not have a significant effect on LLM’s performance for problem-solving tasks [44]. To account for LLM non-determinism, the experiments were executed three times, and the average results were reported, as followed in [45]. The experiments were conducted on a MacBook Pro with an M1 Max chip, 32GB RAM, running macOS Sonoma 14.3.1.
IV-B Research Questions
To evaluate GenLoc, the following research questions are investigated:
RQ1: How does GenLoc compare to other bug localization techniques?
This question analyzes the effectiveness of GenLoc by comparing its performance with existing state-of-the-art techniques using standard evaluation metrics.
RQ2: How do different components of GenLoc contribute to the overall performance?
This question examines the impact of individual components-such as the embedding-based retriever and the LLM reasoning module-on the final bug localization accuracy.
IV-C Baseline, Benchmark Dataset and Evaluation Metrics
GenLoc is compared against a set of techniques with publicly available and workable implementations. We selected five widely studied IRBL techniques that span both traditional and deep learning-based methods: VSM, BugLocator, BLUiR, BRTracer, and DreamLoc. VSM computes cosine similarity between the bug report and source code based on term weights derived using Term Frequency–Inverse Document Frequency (TF-IDF) [46]. BugLocator extends VSM by incorporating file length normalization and historical bug-fix frequency to prioritize more fault-prone files [9]. BLUiR assigns different weights to structural program elements such as class and method names, and computes their similarity to the bug report content separately [47]. BRTracer divides source files into smaller segments (chunks) and integrates additional signals such as stack traces and historical fix data to rank files [29]. DreamLoc leverages a Wide and Deep architecture. The Wide component captures software-specific statistical features (e.g., fix frequency and recency), while the Deep component applies a relevance model for fine-grained similarity scoring [5]. For DreamLoc, the original implementation is used. For BugLocator, BLUiR, and BRTracer, the implementations provided in [1] and for VSM the implementation provided in [46] are adopted since the original implementations are not publicly available. While we aimed to incorporate several other IRBL techniques in the evaluation, they could not be included due to various limitations. For instance, DNNLOC [16] and DeepLoc [17] do not have a publicly released implementation. Although RLocator [21] provides a replication package, it is incomplete due to missing source code. AdaptiveBL [48] could not be executed because of unresolved dependency issues, which also affects FLIM [19], as it builds on AdaptiveBL.
For comparison, the dataset by Ye et al. [20] is used, since it is the most widely used dataset in prior work [5, 13, 48]. It contains over 22,000 bugs from six large, open-source projects, as shown in Table I. Following prior papers [49, 21], 60% data is used as historical (training) data and the most recent 40%—consisting of 9,097 bugs—is used for evaluation.
TABLE I: Dataset Description
Project Time Range # Bug Reports # Java Files LOC
Max Median Min Max Median Min
AspectJ Mar 2002 - Jan 2014 593 6,879 4,439 2,076 699,250 515,153 216,387
Birt Jun 2005 - Dec 2013 4,178 9,697 6,841 1,700 2,322,074 1,549,525 412,795
Eclipse Oct 2001 - Jan 2014 6,495 6,243 3,454 382 1,156,041 637,158 47,348
JDT Oct 2001 - Jan 2014 6,274 10,544 8,184 2,294 922,812 609,378 133,323
SWT Feb 2002 - Jan 2014 4,151 2,795 2,056 1,037 934,357 639,216 301,698
Tomcat Jul 2002 - Jan 2014 1,056 2,042 1,552 924 487,701 413,472 270,046
Similar to other work [9, 50], the following three metrics are used for evaluation:
1. Accuracy@k: It denotes the number of bugs for which at least one of the actual buggy files is ranked within the top-k positions (k = 1, 5, 10) of the ranked list. A higher accuracy@k value indicates a better performance of the bug localization technique. Similar to other studies [5, 19], the value of k is set to 1, 5 and 10.
2. Mean Reciprocal Rank@k (MRR@k): This metric measures the average reciprocal rank of the first correctly identified buggy file within the top-k results across all bug reports. For each bug report, the reciprocal rank is defined as the inverse rank of the first correct buggy file that appears among the top-k results. MRR@k is calculated as follows:
$MRR@k=\frac{1}{N}\sum_{i=1}^{N}\begin{cases}\frac{1}{\text{rank}_{i}}&\text{if }\text{rank}_{i}\leq k\\ 0&\text{otherwise}\end{cases}$ (1)
where, $N$ denotes the total number of bug reports. $\text{rank}_{i}$ refers to the rank of the first correct buggy file for the i-th bug report among the top-k results. If no actual buggy file appears within the top-k results for a bug report, the reciprocal rank is defined as 0. The higher the value of MRR@k, the better the technique. Following existing studies [50, 51], the value of k is set to 10. An MRR@10 value of 0.2 indicates that, on average, the first relevant file is ranked around position 5 within the top 10 results.
3. Mean Average Precision@k (MAP@k): This metric evaluates how well a bug localization technique ranks all actual buggy files within the top-k retrieved results for a set of bug reports. For each bug report, it calculates the Average Precision (AP) using the following formula, which considers both the presence and the ranking positions of all relevant buggy files:
$AP_{i}=\frac{\sum_{j=1}^{M}P(j)\times\text{Rel}(j)}{\text{Number of relevant source files}}$ (2)
where, $M$ is the number of retrieved files in the top-k ranked list for the i-th bug report (i.e., $M$=$k$). $Rel(j)$ denotes whether the file at position $j$ is relevant or not. $P(j)$ indicates the precision at rank $j$ and computed as follows:
$P(j)=\frac{\text{Number of relevant files in top-}j\text{ positions}}{j}$ (3)
The Mean Average Precision is then computed by averaging the AP scores across all bug reports, as shown here:
$MAP=\frac{\sum_{i=1}^{N}AP_{i}}{N}$ (4)
where, $N$ denotes the total number of bug reports. A higher MAP@k value indicates better ranking quality of the buggy files. Following prior studies [50, 51], the value of k is set to 10. For example, a MAP@10 value of 0.3 implies that, on average, 30% of the actual buggy files are retrieved with high precision within the top 10 results.
V Result Analysis
In this section, the experimental results in accordance with the research questions are discussed.
V-A RQ1: Overall Performance of GenLoc
GenLoc is evaluated against five state-of-the-art baseline techniques (DreamLoc, BRTracer, BLUiR, BugLocator, and VSM) across six widely-used benchmark projects (AspectJ, Birt, Eclipse, JDT, SWT, and Tomcat). Tables II and III present the overall and project-wise results, respectively.
TABLE II: Overall Performance of Bug Localization Techniques
Technique Accuracy@k (%) MAP@10 MRR@10
k=1 k=5 k=10
GenLoc 43.19 62.35 67.87 0.41 0.51
DreamLoc 26.33 52.32 63.07 0.30 0.37
BRTracer 26.84 51.21 60.42 0.29 0.37
BLUiR 26.49 49.04 59.82 0.28 0.36
BugLocator 24.19 46.20 56.65 0.26 0.33
VSM 4.01 9.78 14.22 0.05 0.07
V-A1 Overall Results
Table II shows that GenLoc demonstrates strong overall performance across all evaluation metrics compared to prior bug localization techniques. On average, GenLoc achieves the highest Accuracy@1 (43.19%), substantially outperforming the second-best technique (BRTracer with 26.84%) by a relative margin of over 60%. This indicates that GenLoc can correctly identify the buggy file at the top position in a significant portion of bug reports. Similarly, GenLoc leads in Accuracy@5 (62.35%) and Accuracy@10 (67.87%), further affirming its ability to prioritize buggy files effectively within the top candidates. In terms of ranking quality, GenLoc achieves the highest scores in both MAP@10 (0.41) and MRR@10 (0.51), outperforming the next-best MAP@10 (DreamLoc with 0.30) and MRR@10 (BRTracer and DreamLoc with 0.37) by 36.67% and 37.84%, respectively, confirming its superiority in ranking relevant files near the top of the result list.
It is worth mentioning that GenLoc uses only the bug report as input, without incorporating any additional information such as historical bug-fix data or file change frequency. In contrast, techniques like DreamLoc, BRTracer and BugLocator leverage such additional data to improve localization accuracy. Despite considering less information, GenLoc consistently outperforms these baselines on all metrics.
TABLE III: Project-wise Performance Comparison of Bug Localization Techniques
Project Technique Accuracy@k (%) MAP @10 MRR @10
k=1 k=5 k=10
Aspectj GenLoc 46.56 60.48 65.68 0.40 0.53
DreamLoc 14.77 39.66 52.74 0.18 0.26
BRTracer 25.32 55.27 64.56 0.29 0.38
BLUiR 24.05 42.19 58.22 0.23 0.33
BugLocator 18.57 39.24 51.90 0.20 0.28
VSM 2.53 5.06 7.17 0.02 0.04
Birt GenLoc 19.71 34.53 39.88 0.20 0.26
DreamLoc 19.33 39.38 49.13 0.21 0.28
BRTracer 14.00 28.67 37.52 0.14 0.21
BLUiR 13.88 29.80 37.88 0.14 0.20
BugLocator 12.39 26.93 34.95 0.13 0.17
VSM 1.50 4.37 5.75 0.01 0.03
Eclipse GenLoc 46.01 65.99 70.92 0.44 0.54
DreamLoc 34.72 62.36 72.06 0.37 0.46
BRTracer 29.37 54.66 63.74 0.32 0.40
BLUiR 25.52 48.61 57.97 0.28 0.35
BugLocator 28.37 52.12 61.86 0.31 0.38
VSM 5.97 13.32 17.63 0.06 0.09
JDT GenLoc 42.86 66.72 72.92 0.39 0.53
DreamLoc 29.21 54.64 64.48 0.31 0.40
BRTracer 27.62 57.37 64.05 0.28 0.38
BLUiR 30.61 55.08 64.81 0.30 0.41
BugLocator 25.35 51.02 62.38 0.27 0.36
VSM 1.71 4.90 7.85 0.02 0.03
SWT GenLoc 40.96 62.63 70.56 0.41 0.51
DreamLoc 28.92 58.13 69.64 0.34 0.41
BRTracer 19.94 42.35 54.94 0.25 0.30
BLUiR 25.06 51.75 63.98 0.29 0.36
BugLocator 19.94 39.64 51.08 0.23 0.29
VSM 3.13 7.35 11.39 0.04 0.05
Tomcat GenLoc 63.04 83.73 87.28 0.60 0.72
DreamLoc 31.04 59.72 70.38 0.36 0.43
BRTracer 44.79 68.96 77.73 0.46 0.55
BLUiR 39.81 66.82 76.07 0.42 0.51
BugLocator 40.52 68.25 77.73 0.44 0.52
VSM 9.24 23.70 35.55 0.13 0.16
V-A2 Project-wise Results
Table III shows that GenLoc achieves the highest Accuracy@1 across all six projects. Notably, it demonstrates substantial improvements—ranging from 30% to 80%—over the best-performing baseline on projects such as AspectJ, Eclipse, JDT, SWT, and Tomcat. At Accuracy@5, GenLoc continues to outperform all baselines in five out of six projects, most notably achieving 83.73% on Tomcat, 66.72% on JDT, and 65.99% on Eclipse. A similar trend is observed for Accuracy@10, where GenLoc continues to outperform or match other techniques.
Regarding ranking quality, GenLoc achieves the highest MAP@10 and MRR@10 scores on five of the six projects. For instance, on AspectJ, GenLoc achieves an MRR of 0.53, far exceeding that of the second-best technique (BRTracer with 0.38). Similar trends are observed in the MAP scores on projects such as Eclipse, JDT, SWT, and Tomcat, where GenLoc consistently outperforms all baselines, further demonstrating its strength in producing effective ranked outputs.
Bug ID: 324546
Summary: external .js file has no effect in beforeOpen if the data set is used to generate a dynamic parameter selection
Description: Build Identifier: I define a function "replace_query" in an external .js file, and then include it in the report. I then use the function in the beforeOpen script. If the data set is not used to generate a dynamic parameter selection, then it works fine. Otherwise, when viewing the report, it will give this error: Error evaluating Javascript expression. Script engine error: ReferenceError: "replace_query" is not defined. Reproducible: Always Steps to Reproduce: 1. Define a function in an external .js file 2. Include the .js file in the report 3. Call the function in the beforeOpen script of a data set that is also used to generate a dynamic parameter selection.
Figure 3: Bug Report from Birt Project.
Although GenLoc achieves the best Accuracy@1 (19.71%) on Birt project, its Accuracy@5 and Accuracy@10 are lower compared to DreamLoc. Additionally, GenLoc attains the second-highest MAP (0.20) and MRR (0.26), with DreamLoc slightly outperforming it on both metrics. To investigate this performance gap, the authors manually analyzed several Birt bug reports that GenLoc failed to localize and found a common pattern that many of these reports contained “steps to reproduce” information. While such descriptions are helpful for developers [52, 53], they primarily focus on user actions (e.g., clicking buttons or selecting options), which rarely align directly with the source code, as shown in Fig. 3. This disconnection makes it difficult for GenLoc to extract meaningful semantic clues from the report and map them to relevant files.
To systematically verify this observation, reproduction related keywords are extracted from all bug reports following prior work [53]. As shown in Table IV, 63.85% of Birt’s bug reports include such reproduction-related content, which is substantially higher than any other project (e.g., 3.80% in AspectJ, 2.62% in Eclipse, and 2.91% in JDT). This discrepancy contributes to the semantic mismatch between bug reports and source files, which hampers GenLoc’s ability to accurately rank relevant files. Furthermore, a chi-square test revealed a statistically significant relationship between the presence of reproduction information and the failure to localize bugs in the Birt project across all three GenLoc trials (p-value = 0). In contrast, DreamLoc considers additional features such as bug-fixing history of a file, cyclomatic complexity, apart from similarity matching. These supplementary features helped DreamLoc outperform GenLoc on Birt. However, when DreamLoc is executed without these additional features, its performance also drops significantly, as shown in Table LABEL:dreamloc.
TABLE IV: Distribution of bug reports with reproduction-related keywords
Project Total Bugs
Reproduction-related
Keywords
Percentage (%)
AspectJ 237 9 3.80
Birt 1671 1067 63.85
Eclipse 2598 68 2.62
JDT 2509 73 2.91
SWT 1660 88 5.30
Tomcat 422 22 5.21
TABLE V: Comparison of GenLoc and DreamLoc (without additional features) on the Birt Project
Technique Accuracy@k (%) MAP @10 MRR @10
k=1 k=5 k=10
GenLoc 19.71 34.53 39.88 0.20 0.26
DreamLoc (w/o feats) 4.73 11.97 16.16 0.08 0.05
V-A3 Overlap Analysis
Apart from Accuracy@k, MRR@10 and MAP@10, the number of unique and overlapping bugs localized by each approach are examined. In this analysis, a bug is considered successfully localized if the correct buggy file appears within the top 10 ranked results (i.e., Accuracy@10), aligning with a prior study showing that over 98% of developers consider localization results successful when buggy elements are ranked within the top ten positions [54]. Table VI presents the number of total and unique bugs localized by each technique under this criterion. Since GenLoc is a non-deterministic approach, its output varies across different executions due to the inherent randomness in language model sampling. To fairly evaluate GenLoc’s overall capability and compare it with existing deterministic baselines, the union of all bugs localized across multiple GenLoc trials is considered. This union represents the total set of unique bugs that GenLoc is able to localize in at least one run, capturing the full potential of the approach rather than the outcome of a single arbitrary execution.
TABLE VI: Result of Overlap Analysis
Technique Localized Bugs Overlapping Bugs Unique Bugs
GenLoc 6824 6234 590
DreamLoc 5889 5561 328
BRTracer 5283 5250 33
BLUiR 5286 5200 86
BugLocator 5055 5038 17
VSM 1107 1096 11
Among all approaches, GenLoc exhibits the highest overall coverage with 6,824 localized bugs, of which 590 are uniquely identified by GenLoc. This indicates that GenLoc not only localizes more bugs overall but also captures a unique subset of bugs missed by all other methods, highlighting its potential to complement existing methods. This is likely due to its ability to reason over natural language and semantic context, which prior techniques lack.
Bug ID: 55125
Summary: Tomcat does not shut down if LifecycleListener throws an exception
Description: Created attachment 30467 Logs I needed to shut down Tomcat if an application fails to start. I was advised that a LifecycleListener could shut down Tomcat by throwing an IllegalStateException from its lifecycleEvent method. However, this produces severe errors in the logs, but Tomcat does not shut down. I saw this behaviour on Tomcat 7.0.41 and 7.0.40 using the test web application from here: https://github.com/cloudfoundry/java-test-applications and v1.0.0 of the listener from here: https://github.com/cloudfoundry/java-buildpack-support (v1.1.0 of the listener was re-coded to halt the JVM). I attach catalina.out. Mark Thomas suggested I raise this bug, so he believes that Tomcat should shut down.
Figure 4: Bug Report from Tomcat Project.
V-A4 Case Study
A sample bug [^2] https://bz.apache.org/bugzilla/show_bug.cgi?id=55125 from Tomcat project is shown in Fig. 4, where all prior techniques such as BugLocator, BLUiR, and DreamLoc failed to localize the bug due to vocabulary gap between the bug report and source code. In contrast, GenLoc could place the correct buggy file Catalina.java at the top (rank 1). It first understood that the core issue in the bug report was related to Tomcat failing to shut down when a LifecycleListener throws an exception. Based on this understanding, it identified files related to lifecycle events by invoking get_candidate_filenames() and search_file(). It then used get_method_signatures_of_a_file() to inspect the structure of these files and locate shutdown-related methods. Finally, by analyzing the method bodies using get_method_body(), it examined functions like start() and stop() in Catalina.java, ultimately pinpointing it as the buggy file. This example highlights GenLoc’s strength in reasoning about semantic intent and program behavior, going beyond surface-level text matching.
V-B RQ2: Ablation Study
To better understand the contribution of different components within GenLoc’s bug localization pipeline, an ablation study is conducted. In particular, the goal is to examine how embedding-based file retrieval and LLM-based iterative reasoning interact to improve bug localization performance. By selectively disabling or isolating these components, their individual and combined impact are examined. For this purpose, the following three configurations are used:
1. GenLoc: This is the complete version of the proposed approach, where the LLM is provided with five functions, including get_candidate_filenames() to utilize embedding-recommended files.
2. Embedding-Only: This variant evaluates the standalone performance of the embedding-based retrieval system, where source files are ranked solely based on their semantic similarity to the bug report, without any reasoning or iteration by the LLM.
3. GenLoc-NoEmbed: In this configuration, the LLM operates without access to the embedding-based recommended files, i.e., the get_candidate_filenames() function is removed. The model must infer filenames directly from the bug report and continue reasoning with the remaining four functions.
Table VII presents the average results across all projects. GenLoc achieves the best performance on all metrics, demonstrating the strength of combining embedding-based retrieval with iterative LLM analysis. The Embedding-Only variant obtains only 32.59% at Accuracy@1, indicating that retrieval alone is insufficient without deeper reasoning. A sample case [^3] https://bugs.eclipse.org/bugs/show_bug.cgi?id=232757 is shown in Fig. 5, where Embedding-Only was unable to localize the correct file, while GenLoc succeeded. Due to space constraints, a selection of representative lines from the bug report is presented. The Embedding-Only approach failed to localize the correct buggy file due to excessive noise in the bug report, which mentioned over 50 unique filenames, most of which were unrelated to the fault. GenLoc overcame this challenge by first identifying key method names from the stack trace (e.g., zoomOut() and render()), then searching for files containing these methods. It subsequently examined the method bodies to assess their relevance to the failure. This structured reasoning enabled GenLoc to correctly localize the buggy file, AutoScale.java, despite the presence of substantial noise in the report.
TABLE VII: Result of Ablation Study
Technique Accuracy@k (%) MAP@10 MRR@10
k=1 k=5 k=10
GenLoc 43.19 62.35 67.87 0.41 0.51
Embedding-Only 32.42 56.86 65.50 0.34 0.43
GenLoc-NoEmbed 34.28 43.29 45.71 0.31 0.38
Bug ID: 232757
Summary: Error with standard meter charts: org.eclipse.birt.chart.exception.ChartException: Could not zoomOut for step=8
Description: I am using BIRT 2.2.2. I have a report that displays several “standard meter” charts. But in some cases, some dials are not displayed and I get the following error:
org.eclipse.birt.chart.exception.ChartException: Could not zoomOut for step=8
at org.eclipse.birt.chart.factory.Generator.render(Generator.java:1287)
org.eclipse.birt.report.engine.layout.html.HTMLPageLM.layout(HTMLPageLM.java:85) org.eclipse.birt.chart.computation.withaxes.AutoScale.zoomOut(AutoScale.java:515) … 151 more. Also I have put in attachment a simplified version of a report that shows problems.
Figure 5: Bug Report from Birt Project.
The GenLoc-NoEmbed achieves only 34.61% Accuracy@1, showing that the absence of embedding guidance limits the LLM’s ability to identify relevant files effectively. Fig. 6 presents a sample bug [^4] https://bugs.eclipse.org/bugs/show_bug.cgi?id=147801 from the AspectJ project, where GenLoc successfully localized the bug, while GenLoc-NoEmbed failed. The bug report referenced a DB2 class (DB2PreparedStatement) and indicated a failure during runtime weaving [^5] Runtime weaving denotes the process of dynamically inserting additional behavior, known as aspects (e.g., logging), into a program during execution, without modifying the original source code.. Although both GenLoc and GenLoc-NoEmbed attempted to find the mentioned class or methods directly, these efforts were unsuccessful. However, GenLoc overcame this problem by invoking get_candidate_filenames(), which retrieved semantically related weaving components such as BcelClassWeaver.java and WeavingAdaptor.java. Next, by analyzing their method bodies like weave() and weaveClass(), GenLoc could identify the correct buggy file (BcelClassWeaver.java)). In contrast, GenLoc-NoEmbed lacked access to these relevant files due to its inability to retrieve semantically similar candidates, demonstrating the importance of embedding-guided retrieval in resolving non-obvious fault locations. Overall, these results highlight that embedding-based retrieval and LLM-based reasoning are complementary, and removing either degrades localization performance.
Bug ID: 147801
Summary: java.lang.ClassFormatError: Repetitive method name/signature
Description:I get a ClassFormatError trying to perform runtime-weaving on a DB2 class. The class being woven is COM.ibm.db2.jdbc.app.DB2PreparedStatement. The full message exception is: java.lang.ClassFormatError: Repetitive method name/signature in class file COM/ibm/db2/jdbc/app/DB2PreparedStatement. In examining the post-weave class file, it looks as if Aspectj is defining a duplicate method within the class called getParameterMetaData().
’COM.ibm.db2.jdbc.app.DB2PreparedStatement’ (DB2PreparedStatement.java:3971) advised by afterReturning advice from ’com.ibm.tivoli.itcam.toolkit.ai.aspectj.captureJDBC.CaptureSQLStatement’ (CaptureSQLStatement.aj:946) [with runtime test]
Figure 6: Bug Report from AspectJ Project.
VI Discussion
This paper presents the first IR-based bug localization technique that uses an LLM. As highlighted in a recent survey  [55], future research should focus on leveraging state-of-the-art LLMs such as GPT and LLaMA and exploring prompt engineering and agent-based techniques to enhance IR-based bug localization. GenLoc operationalizes this vision by integrating semantic retrieval and LLM-driven reasoning within a unified bug localization pipeline.
The results demonstrate that GenLoc outperforms existing bug localization techniques, achieving the highest Accuracy@1 and consistently ranking buggy files higher across six diverse open-source Java projects. Compared to traditional IR-based techniques such as BugLocator and BRTracer, as well as recent deep learning-based methods like DreamLoc, GenLoc delivers substantial improvements by combining embedding-based candidate retrieval with iterative reasoning powered by an LLM. Unlike BugLocator, BRTracer, and DreamLoc, GenLoc operates solely on the bug report, without relying on historical bug-fix data or change history—resources that are often unavailable in new projects. This makes GenLoc more broadly applicable to any project.
The average cost to localize a bug using GenLoc is \$0.012, which includes \$0.003 for embedding-based retrieval (Step-1) and \$0.008 for LLM-based reasoning (Step-2). The average processing time per bug is 48.40 seconds (8.32 seconds for Step-1 and 40.08 seconds for Step-2). Notably, this processing time aligns well with developer expectations where prior studies indicate that 90% of developers expect a bug localization technique to return results within one minute. GenLoc meets this requirement while maintaining high accuracy and low cost, making it practical for real-world debugging scenarios.
Despite its success, GenLoc has several limitations. It may fail to localize bugs when the bug report lacks semantic or behavioral clues—particularly when the report is written from a user perspective and focuses mainly on reproduction steps without mentioning relevant program behavior or entities. One promising direction to address this challenge is to integrate an LLM-driven query reformulation step that rewrites or enriches the original bug report to better align with the source code and improve the quality of retrieved candidates.
Another limitation is that GenLoc may not identify all files associated with a given bug within its top-10 ranked list, which is the current output size of the system. However, a prior work from Facebook [56] suggests that predicting even a single relevant file is useful in practice since it helps route the issue to the appropriate developer or team and can serve as a starting point for discovering additional affected files through manual inspection or complementary techniques. Although this does not replace the need for comprehensive localization, it highlights the potential value of partial predictions in real-world debugging workflows.
VII Threats to Validity
This section presents potential ways in which the validity of the study may be compromised.
Internal Validity: One of the most significant threats to internal validity is the possibility of data leakage, such as bug-fixing commits being present in the LLM’s training data. However, the comparison between GenLoc and GenLoc-NoEmbed in RQ2 suggests that GenLoc’s performance is not solely attributable to model memorization. Another potential threat arises from the inherent randomness of LLMs, which may yield different results across different runs. To mitigate this, all experiments were repeated three times, and a replication package is released to support reproducibility and enable further validation by the research community. Additionally, the choice of LLM may impact the results, as different models can vary significantly in reasoning capabilities, code understanding and sensitivity to prompt structure. Hence, future evaluations should assess the robustness of GenLoc across multiple LLM variants.
Construct Validity: Construct validity concerns whether the evaluation metrics accurately measure the effectiveness of bug localization techniques. This study employs Accuracy@k, MRR@10, and MAP@10, which are widely used in prior bug localization research and considered standard for assessing ranking quality [2].
External Validity: GenLoc is evaluated on over 9,000 bug reports from six open-source Java projects of varying sizes and domains. This dataset is widely used in prior bug localization research [21]. However, the findings may still not generalize to proprietary industrial systems or projects written in other programming languages. Additional studies are needed to evaluate the effectiveness of GenLoc in broader and more heterogeneous settings.
VIII Related Work
This section discusses papers from two related domains.
VIII-A Information Retrieval-based Bug Localization
Existing IRBL approaches can be broadly categorized into two groups based on their underlying methodology. The first category [57, 12, 9, 58, 59, 47, 29] uses traditional IR models such as Latent Dirichlet Allocation (LDA) and VSM for bug localization. While Lukins et al. were the first to apply LDA for bug localization [57], BugScout extended this idea by adding bug history and file size [12]. BugLocator introduced a revised VSM incorporating file size and similar past bugs [9]. BLUiR further improved retrieval by using structured code elements such as class and method names [47], while AmalGAM combined the strengths of BugLocator and BLUiR with version history [60]. Sisman and Kak introduced bug and file modification history [58], and Kim et al.improved precision by filtering uninformative bug reports [59]. BRTracer included stack traces and code chunking [29], while later efforts such as Locus [61] and BLIA [62] incorporated fine-grained change information.
The second category [63, 48, 15, 16, 17, 64, 5, 65, 66, 19, 67, 68, 21, 69, 70, 71] includes machine learning and deep learning based approaches. [63, 48] leveraged learning-to-rank model for bug localization. Early deep learning methods such as HyLoc [15], DNNLOC [16], and DeepLoc [17] used DNNs to find semantic similarity between bug report and source code. CNN-based models like BugPecker [64], DreamLoc [5], MD-CNN [65], and TRANP-CNN [66] enhanced structural understanding using ASTs, CFGs, and transfer learning. Recent approaches, including FLIM [19], FBL-BERT [67], HMCBL [68] and RLocator [21] leveraged pre-trained language models and hybrid architectures to improve semantic matching and scalability. Other models like CoLoc [69], ImbalancedBugLoc [70], and DependLoc [71] addressed specific challenges such as class imbalance, cross-project localization, and dependency-aware analysis. Unlike other approaches, this paper explores the use of large language models for IRBL.
VIII-B LLMs in Software Debugging and Automated Issue Resolution
Recent studies have investigated the usage of LLMs in software debugging and automated issue resolution. A notable line of work focuses on spectrum-based bug localization, where approaches such as AutoFL [72], LLM4FL [73], and CosFL [74] leverage failing test cases to identify faulty code segments. Building on this, FuseFL [75] enhances interpretability by generating step-by-step reasoning to justify faulty lines, while FlexFL [76] further improves accuracy by combining bug reports with execution traces.
In parallel, other research efforts aim to enhance bug reports using LLMs. For instance, ChatBR [77] generates missing information such as observed and expected behavior, and Plein and Bissyandé [78] investigate how LLMs can generate test cases directly from bug reports. However, none of these studies analyzed the impact of LLM in the context of IRBL, when no other information (e.g. test cases) is available.
Another emerging direction involves the use of LLMs for automated issue resolution on platforms like GitHub. Techniques such as MAGIS [79], AutoCodeRover [45], SpecRover [80] and Agentless [81] aim to address software issues including bugs. However, these methods primarily focus on the repair phase, often overlooking the bug localization step, which remains a crucial and challenging part of the debugging process [82].
IX Conclusion and Future Work
This paper presents GenLoc, a novel information retrieval-based bug localization technique that integrates semantic retrieval with step-by-step reasoning capabilities of LLMs. An evaluation on over 9,000 real-world bugs from six diverse Java projects shows that GenLoc outperforms existing techniques in terms of Accuracy@k, MRR@10, and MAP@10. The results highlight GenLoc’s ability to localize bugs that lack direct lexical overlap with code, a major challenge for conventional IR-based methods. Moreover, GenLoc can localize the highest number of unique bugs, demonstrating its potential to complement existing approaches.
In future, GenLoc could be extended to support method-level bug localization to provide finer-grained debugging assistance. Furthermore, the impact of different LLMs on both the effectiveness and efficiency of localization could be analyzed.