### v2

# Leveraging Large Language Model for Information Retrieval-based Bug Localization

Abstract.
Information Retrieval-based Bug Localization (IRBL) aims to identify buggy source files for a given bug report. Traditional and deep-learning-based IRBL techniques often suffer from vocabulary mismatch and dependence on project-specific metadata, while recent Large Language Model (LLM)-based approaches are limited by insufficient contextual information. To address these issues, we propose GenLoc, an LLM-based technique that combines semantic retrieval with code-exploration functions to iteratively analyze the code base and identify potential buggy files. We evaluate GenLoc on two diverse datasets: a benchmark of 9,097 bugs from six large open-source projects and the GHRB (GitHub Recent Bugs) dataset of 131 recent bugs across 16 projects. Results demonstrate that GenLoc substantially outperforms traditional IRBL, deep learning approaches and recent LLM-based methods, while also localizing bugs that other techniques fail to detect.
bug localization, debugging, large language model, information-retrieval
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Software and its engineering Software maintenance toolsccs: Information systems Information retrieval
1. Introduction
Bug localization is the process of identifying the faulty or buggy locations in source code (Lee et al., 2018). It is a crucial step in software debugging to ensure the quality of a software (Wong and Tse, 2023). However, localizing bugs in practice is challenging, as large systems often incur a huge number of bugs everyday. For example, Mozilla receives around 300 new bug reports daily, which is overwhelming for the developers (Anvik et al., 2006). Moreover, given the size and complexity of large-scale software systems today, manually locating bugs is a time-consuming, tedious and expensive task (Wong et al., 2016). Consequently, there is a strong demand for automated bug localization techniques to ease the burden of developers and accelerate the program debugging process (Qi et al., 2021).
Over the years, researchers have proposed multiple approaches for bug localization. Among these, Spectrum-based Bug Localization (SBBL) and Information Retrieval-based Bug Localization (IRBL) are the most studied and widely recognized techniques (Liu et al., 2024; Qi et al., 2021). SBBL leverages runtime behavior (e.g., test coverage information) to analyze which parts of a program are exercised by passing and failing test cases, and ranks program elements based on their likelihood of being faulty (Zou et al., 2019). In contrast, IRBL takes a bug report—a document that describes a software bug—as input and does not require bug-revealing test cases or execution traces (Kim and Lee, 2021). Due to its simplicity, IR-based approaches have received growing attention in recent years.
IRBL approaches consider the bug report as a query and the software repository as a collection of documents (i.e., source files). Given a bug report, these approaches output a ranked list of source files based on their probability of being faulty. The goal is to rank the actual buggy files as high as possible. To calculate the similarity between bug reports and the source files, earlier techniques use Vector Space Model (VSM) (Zhou et al., 2012; Thomas et al., 2013), Topic Modeling (Lukins et al., 2010; Nguyen et al., 2011), while later methods (Ye et al., 2016; Liu et al., 2019; Lam et al., 2015, 2017; Xiao et al., 2019; Zhang et al., 2020; Qi et al., 2021; Liang et al., 2022; Ye et al., 2015; Chakraborty et al., 2024) rely on machine learning and deep learning models. Many of these approaches also incorporate additional information such as historical bug-fix data. However, such information is not always available—particularly in new projects—thereby limiting the generalizability of these methods. Furthermore, the effectiveness of these approaches is often limited by the vocabulary gap between bug reports and source code, as they tend to use different terminologies (Liu et al., 2024).
To bridge these limitations, recent studies have incorporated Large Language Models (LLMs) into IRBL, as they can reason over both natural language and source code (Samir and Rahman, 2025; Li et al., 2025a). These approaches generally retrieve a small set of candidate files and then use an LLM for further analysis or reranking. For instance, BRaIN shortlists files using lexical similarity (Samir and Rahman, 2025), whereas LLM-BL retrieves candidates using hybrid search (Li et al., 2025a). While such methods improve upon traditional IRBL techniques, their reliance on a restricted context (e.g., a small set of shortlisted files) can result in faulty files being overlooked. If the actual buggy file is absent from the retrieved candidates, the LLM will fail to localize the bug. Expanding the context is a natural alternative, but it introduces new challenges. For example, providing the entire repository as input is infeasible due to context window constraints, and examining files one by one is impractical for large projects. Furthermore, LLMs often struggle to identify relevant information when the input context is long and relevant information is buried in the middle (Liu et al., 2023). These challenges highlight the need for an effective strategy that enables LLMs to navigate the code base, selectively retrieve relevant information and reason within a constrained context window.
In this context, we propose a novel bug localization technique, named GenLoc, that combines semantic retrieval with iterative reasoning capabilities of LLMs to identify buggy files for a given bug report. The name GenLoc stands for Generative AI-based Bug Localization, emphasizing its core design principle of leveraging generative language models for bug localization. Unlike many prior approaches (Zhou et al., 2012; Qi et al., 2021), GenLoc relies solely on the bug report and does not require any additional project-specific metadata or historical bug-fix data. At first, GenLoc converts the source code into chunks and measures their semantic similarity with the bug report to retrieve a set of potentially relevant files. Next, it asks an LLM to analyze a bug report and inspect these retrieved files further, or explore other files as needed. To support this exploration, the LLM is equipped with a set of functions (e.g., retrieving method signatures or method bodies) that enable it to iteratively reason about which files are likely related to the reported bug. This design enables guided and selective localization, avoiding the pitfalls of narrow retrieval that excludes the faulty file and excessive repository-level input that overwhelms the LLM.
We evaluate GenLoc on two complementary datasets: (i) a large benchmark of 9,097 bug reports from six open-source projects that has been widely used in prior bug localization research (Chakraborty et al., 2024), and (ii) the more recent GHRB (GitHub Recent Bugs) dataset of 131 bugs from 16 projects, designed specifically for assessing LLM-based approaches (Lee et al., 2024). Results show that GenLoc consistently outperforms existing approaches. On the large benchmark, GenLoc improves Accuracy@1 by over 63% compared to traditional and deep learning-based IRBL techniques, and on the GHRB dataset, it surpasses recent LLM-based methods by more than 25%. GenLoc also achieves superior ranking performance by placing correct files much earlier in the result list. Moreover, it exhibits complementary strengths to existing techniques, underscoring its potential for integration into hybrid approaches to advance bug localization performance further.
In summary, this paper makes the following contributions:
  • The introduction of a novel bug localization approach named GenLoc that integrates semantic retrieval with LLM-guided iterative reasoning.
  • A comprehensive empirical evaluation of GenLoc on more than 9,000 real-world bugs from six large open-source projects and the recent GHRB dataset, including comparisons with traditional IRBL techniques and state-of-the-art LLM-based approaches, as well as an ablation study to assess the contribution of each component.
  • A publicly accessible replication package to facilitate future research.
2. Background
This section outlines the key technical components that form the foundation of the proposed approach.
Embedding: It represents complex data, such as text, images, or source code, as numerical vectors in a high-dimensional space. The core idea is that semantically similar items (e.g., two similar sentences or code snippets) will have embeddings that are close to each other in that space. In GenLoc, both source files and bug reports are converted into embeddings to enable semantic similarity calculation between them.
Vector database: A vector database is a special type of database designed to store and manage large collections of high-dimensional embeddings and perform efficient similarity searches (Han et al., 2023b). To accelerate the retrieval process, vector databases commonly use Approximate Nearest Neighbor (ANN) algorithms (Tian et al., 2023), which find vectors that are close to a given query without exhaustively comparing all entries. In GenLoc, the embeddings of all the source files are stored in a vector database to facilitate retrieval of the top source files that are most semantically similar to a given bug report.
Function calling: This mechanism allows LLMs to interact with external tools and systems to retrieve up-to-date information or perform specific tasks that they cannot directly do. For example, if LLM is given the prompt "What is the current temperature in Los Angeles?", it will fail to answer because it lacks access to real-time data. However, it can be provided with access to a function like get_current_temperature(), which queries an external weather API and returns the latest temperature. The model can then call this function, receive the temperature and integrate it into its response. In GenLoc, the LLM is given access to several functions such as retrieving method signatures and bodies that allow it to iteratively navigate the code base and identify potentially buggy files.
ReAct framework: The ReAct (Reasoning and Acting) framework enables language models to interleave natural language reasoning with tool-based actions, thereby enhancing their problem-solving capabilities (Yao et al., 2023). Instead of generating a final answer in a single step, the model alternates between articulating intermediate reasoning steps and performing actions, such as calling a function. This synergy between reasoning and acting allows the model to iteratively refine its understanding, validate hypotheses, and make informed decisions. Since ReAct has shown promise in handling complex tasks that require multi-step thinking, access to external environments, or dynamic decision-making (Yao et al., 2023; Wu, 2025), this paper adopts the ReAct framework to guide the LLM in navigating the code base and progressively narrowing down the search space during bug localization.
Figure 1. Workflow of GenLoc.
3. Methodology
GenLoc operates in two primary steps to localize relevant files based on a given bug report. First, it retrieves a set of semantically similar files using embedding-based similarity. Next, an LLM, supported by a set of external functions, iteratively analyzes the bug report and the code base. During this stage, the model may examine the embedding-based retrieved files or explore other parts of the code base to generate a ranked list of potentially buggy files. An overview of the workflow is shown in Fig. 1.
3.1. Shortlisting suspicious files
To identify source files potentially related to a given bug report, this step computes the semantic similarity between the bug report and the source code using embedding-based representations. The process begins by traversing all source files in the project. For each file, its fully qualified file path and method bodies are extracted. Each file is then represented as a concatenation of its fully qualified file path and the method bodies. The inclusion of the file path provides contextual information about the file’s role or feature within the overall system architecture (e.g., org.eclipse.jdt.ui/ui/org/eclipse/jdt/ui/JavaElementLabels.java suggests a UI-related component), while method bodies capture the actual implementation logic, which is crucial since methods serve as the fundamental units of behavior in a program.
The concatenation of a file’s path and method bodies can result in lengthy text. To better localize relevant content, this representation is split into smaller chunks if it exceeds a predefined length threshold (Wong et al., 2014). Each chunk is then converted into an embedding vector and stored in a vector database. Similarly, the textual content of the bug report—comprising its summary and description—is transformed into embedding vector.
To measure the similarity between the bug report and the source code, cosine similarity is used, which measures the angular distance between the embedding vectors (Steck et al., 2024). Based on the similarity scores, a predefined number of the top suspicious files are shortlisted. These files serve as candidate inputs for further analysis by the LLM in the subsequent step.
3.2. Ranking suspicious files
In this step, an LLM is employed to generate a ranked list of potentially buggy source files based on a given bug report, as illustrated in Fig. 1. To enhance the LLM’s reasoning capabilities, a prompt (shown in Fig. 2) is constructed by following the ReAct (Reasoning and Acting) framework (Yao et al., 2023), which enables LLMs to interleave natural language reasoning with tool-based actions. This approach has proven effective in improving performance on complex decision-making tasks by enabling iterative reasoning and environment interaction.
You are an expert software engineer specializing in bug localization. Your goal is to identify the most probable buggy Java files based on a given bug report. You have access to five functions that help you search for file names, locate methods and analyze source code. You must follow an iterative, reasoning-based approach, refining your strategy dynamically based on insights gathered during each step. At each iteration, you should: Maintain a working shortlist of files that appear potentially buggy. Update the shortlist based on new evidence (e.g., method matches, code analysis). Avoid redundant operations–do not recheck the same filenames, methods or file contents multiple times. Continue this process until you either: (a) Produce a well-justified ranked list of the 10 most relevant files, or (b) Reach the maximum limit of 10 iterations. In the 10th iteration, you must output your final ranked list. Workflow 1. Analyze the Bug Report: Extract relevant keywords, error messages and functional hints from the bug summary and description. Identify potentially affected components (e.g., UI, database, networking). 2. File Discovery: Use search_file() to check if filenames derived from the bug report’s keywords or functionality exist in the code base. If the bug report references a specific method name, use search_method() to find all files defining it. If an inferred filename or method location does not exist, refine your strategy: adjust assumptions, explore variations and retry. If strong matches are not found, use get_candidate_filenames() to retrieve 50 potentially relevant filepaths. Add promising files to your shortlist by prioritizing those align with terminology, functionality or methods discussed in the bug report. 3. Method Analysis: For each file in the shortlist that hasn’t been analyzed yet: Use get_method_signatures_of_a_file() to list its methods. Identify methods that directly align with the bug’s context (e.g., related functionality, naming hints). For any method of interest, retrieve its implementation using get_method_body(). Analyze the logic of any identified method(s) of interest to determine whether they align with the bug’s symptoms. If this analysis reveals new relevant class names, filenames, or methods, use the appropriate search functions. Continuously update the shortlist by promoting, demoting or removing files based on evolving understanding. 4. Shortlist Refinement & Ranking: Rank files based on: Semantic alignment with the bug report’s keywords and described functionality Method or filename alignment with bug context Code logic alignment with the bug description If needed, iterate with refined assumptions or explore previously overlooked filenames or methods. 5. Final Output: Provide a ranked list of the 10 most relevant filepaths based on their likelihood of containing the bug. Ensure filepaths exactly match those provided–do not modify case, structure or abbreviate them. Justify each file’s inclusion by referencing keywords, method matches or code logic that supports its relevance to the bug report.
Figure 2. LLM Prompt.
At first, the LLM is prompted to assume the role of an expert software engineer specializing in bug localization, as role-playing improves the reasoning abilities of LLMs (Kong et al., 2023). Next, the prompt is designed to decompose bug localization into a series of smaller tasks since LLMs perform better when complex tasks are broken down into sub-tasks (Khot et al., 2022). These sub-tasks include analyzing the bug report, identifying potential file or method names, analyzing method signatures and bodies and integrating evidence to prioritize suspicious files.
To support the reasoning process, the LLM is given access to the following five external functions. The LLM may invoke one or more of these functions over a fixed number of iterations to get information about the code base. Collectively, these functions allow the LLM to discover files, understand their role and analyze the internal logic of specific code segments, thereby supporting both breadth and depth in the bug localization process.
  1. (1)
    search_file(): This function enables the LLM to check whether a specific file exists in the code base. It can be used to (i) verify the existence of a file explicitly mentioned in the bug report (e.g., checking if JavaElementLabels.java exists when “JavaElementLabels” is referenced), (ii) infer and validate a possible filename derived from textual clues in the bug report, or (iii) identify candidate files after inspecting a method body.
  2. (2)
    search_method(): When a method name is mentioned in the bug report, e.g., updateLabel(), this function enables the LLM to find which files contain its definition and thereby narrowing down the search space. In addition, it can be employed to discover candidate files based on references observed during method body inspection.
  3. (3)
    get_candidate_filenames(): This function retrieves a list of potentially relevant files based on their semantic similarity to the bug report from the previous step (Section 3.1). The LLM can leverage this list as an initial pool of candidates, especially when the bug report does not provide any clues.
  4. (4)
    get_method_signatures_of_a_file(): Once a file is identified, this function enables the model to inspect the file’s high-level structure by retrieving all method signatures defined within it. This helps the LLM determine whether the file’s responsibilities align with the bug report.
  5. (5)
    get_method_body(): To evaluate a method’s implementation, the model can call this function to retrieve the body of a given method. This allows a deeper inspection of whether the method may be related to the described issue.
At each iteration, the model can “reason” about what information is currently available, determine what additional data is required, and then “act” by invoking one of these functions. The outcome of the function call is integrated into the LLM’s internal reasoning in the subsequent iteration. Since the LLM can make mistakes while invoking function calls, such as supplying incorrect parameters like non-existent filenames or method signatures, we incorporate an iteration-time recovery step. When no exact match is found, the system either provides tentative options (e.g., all file paths sharing the same base filename or the closest method signatures using Damerau–Levenshtein distance (Zhao and Sahni, 2019)) or returns explicit feedback indicating that the requested element does not exist in the code base. This mechanism guides the LLM in handling mistakes (e.g, typos or incomplete names) by recovering the intended file or method, or preventing it from producing non-existent ones.
Lastly, the LLM is prompted to output a ranked list where each entry includes a fully qualified file path along with a brief justification for its selection. Since multiple files from different packages can have the same name, the fully qualified file path is used to uniquely identify each file. The inclusion of justifications is motivated by prior work demonstrating that self-explanatory prompts significantly improve LLM comprehension (Wu et al., 2023).
Although the LLM receives feedback during intermediate iterations, it can still produce erroneous or non-existent file paths in the final ranked list. A common issue we observed in our preliminary analysis is the omission of intermediate packages. For example, the model may omit the "internal" package and produce org.eclipse.jdt.ui/ui/org/eclipse/jdt/ui/text/java/AlphabeticSorter.java instead of org.eclipse.jdt.ui/ui/org/eclipse/jdt/internal/ui/text/java/AlphabeticSorter.java. To address this, we apply a post-processing recovery step to verify the correctness of each predicted file. If the predicted fully qualified file path matches an existing file, it is retained. Otherwise, the base filename (e.g., AlphabeticSorter.java) is extracted, and all files with the same name are retrieved. Among these candidates, the file whose fully qualified file path has the highest Jaccard similarity (Niwattanakul et al., 2013) with the predicted filename is selected. Since Jaccard measures token overlap, it favors the correct file even when an intermediate package is missing. If the base filename does not exist in the code base, no suitable match can be found and the file is excluded from the final ranked list.
4. Experiement Setup
This section presents the implementation details, research questions, baseline techniques, benchmark dataset and evaluation metrics.
4.1. Implementation
Similar to most of the existing techniques (Ye et al., 2015; Qi et al., 2021; Chakraborty et al., 2024), GenLoc focuses on Java projects since it is one of the most popular programming languages. However, the approach is easily extendable to other languages, as it leverages the Tree-sitter library (tre, [n. d.]) for source code parsing, which supports more than ten programming languages. To compute semantic similarity between bug reports and source code files, GenLoc employs OpenAI’s text-embedding-3-small model due to its cost-effectiveness (\$0.02 per 1M tokens) (tex, [n. d.]). Following Amazon Bedrock’s recommendation (aws, [n. d.]), we use a chunk size of 300 tokens to balance between granularity and context—capturing relevant code segments while preserving surrounding information for meaningful semantic comparison. For each bug report, the top 50 source files are retrieved based on their embedding similarity to the report (Li et al., 2025a).
To improve efficiency and minimize embedding costs, embeddings for all source files are generated and stored in ChromaDB (chr, [n. d.]) when processing the first bug in each project. Since different bugs occur in different versions, for each subsequent bug, only the embeddings of files that have been added, modified, deleted, or renamed across versions are updated in the ChromaDB. For inference, GPT-4o mini by OpenAI is used due to its cost-effectiveness (\$0.15 per 1M input tokens and \$0.60 per 1M output tokens) (gpt, [n. d.]). All model parameters are kept at their default settings to avoid the computational cost of hyperparameter tuning. In particular, the temperature parameter is left at its default value (1.0) (rather than being set to 0) to enable diverse reasoning paths during inference (Kang et al., 2023; Zhu et al., 2024). A temperature of 0 significantly reduces variability by consistently selecting the highest-probability tokens, which can lead the model to overlook less obvious but relevant files. Our preliminary analysis also showed that the performance remained mostly unchanged regardless of the temperature setting. This aligns with prior findings that changes in temperature between 0.0 and 1.0 do not have a significant effect on LLM’s performance for problem-solving tasks (Renze, 2024). To account for LLM non-determinism, the experiments were executed three times and the average results were reported, as followed in (Zhang et al., 2024).
4.2. Research Questions
To evaluate GenLoc, the following research questions are investigated:
  • RQ1: How does GenLoc perform compared to traditional and deep learning-based IRBL techniques?
  • RQ2: How does GenLoc perform compared to other LLM-based IRBL techniques?
  • RQ3: How do different components of GenLoc contribute to the overall performance?
  • RQ4: How well does GenLoc perform on previously unseen bugs?
4.3. Baseline Techniques
To provide a comprehensive evaluation of GenLoc, we compare it against a diverse set of file-level IRBL techniques with publicly available and executable implementations. The baselines cover both traditional, deep learning-based methods and recent LLM-based approaches.
We selected four widely studied IRBL techniques from the traditional and deep learning categories: BugLocator (Zhou et al., 2012), BLUiR (Saha et al., 2013), BRTracer (Wong et al., 2014) and DreamLoc (Qi et al., 2021). BugLocator uses revised Vector Space Model (rVSM) by incorporating file length normalization and historical bug-fix frequency to prioritize more fault-prone files (Zhou et al., 2012). BLUiR assigns different weights to structural program elements such as class and method names, and computes their similarity to the bug report content separately (Saha et al., 2013). BRTracer divides source files into smaller segments (chunks) and integrates additional signals such as stack traces and historical fix data to rank files (Wong et al., 2014). DreamLoc leverages a Wide and Deep architecture. The Wide component captures software-specific statistical features (e.g., fix frequency and recency), while the Deep component applies a relevance model for fine-grained similarity scoring (Qi et al., 2021).
For DreamLoc, the original implementation is used. For BugLocator, BLUiR and BRTracer, the implementations provided in (Lee et al., 2018) are adopted, since the original implementations are unavailable. While we aimed to incorporate several other IRBL techniques in the evaluation, they could not be included due to implementation-related issues. For instance, DNNLOC (Lam et al., 2017) and DeepLoc (Xiao et al., 2019) do not have a publicly released implementation. Although RLocator (Chakraborty et al., 2024) provides a replication package, it is incomplete due to missing source code. AdaptiveBL (Fejzer et al., 2021) and its successor FLIM (Liang et al., 2022) could not be executed due to unresolved dependency issues.
In addition, we compare GenLoc against three LLM-based approaches: BRaIn (Samir and Rahman, 2025), LLM-BL (Li et al., 2025a) and Agentless (Xia et al., 2025). BRaIn retrieves top-K files using lexical search, identifies relevant methods from those files using an LLM and refines the results through keyword-based query expansion and reranking (Samir and Rahman, 2025). LLM-BL expands the bug report using an LLM, retrieves candidate files through hybrid (lexical and semantic) search and presents the shortlisted files along with their relevant code lines to the LLM for reranking (Li et al., 2025a). Agentless is the leading open-source issue-fixing technique and has been adopted by OpenAI as the standard approach for showcasing real-world coding performance (Xia et al., 2025; Chang et al., 2025). It localizes bugs through three steps: (i) representing the code base as a tree-like structure and prompting the LLM to rank suspicious files, (ii) using embeddings to retrieve additional candidates, and (iii) combining both results into the final list of suspicious files. For both BRaIn and LLM-BL, we use their original implementations, while for Agentless, we extract its file-level bug localization component and adapt it (e.g., code parsing) for Java projects, since the original version only supports Python.
4.4. Benchmark Dataset
To evaluate GenLoc, we employ two complementary benchmark datasets: the widely used Ye et al. dataset (Ye et al., 2015) and the more recent GHRB dataset (Lee et al., 2024). For RQ1, we use the Ye et al. dataset which is the most widely adopted benchmark in prior IRBL studies (Qi et al., 2021; Ye et al., 2016; Fejzer et al., 2021). This dataset includes over 22,000 bugs from six large, open-source projects (Table 1), which makes it suitable for training deep-learning based approach DreamLoc. Following prior work (Han et al., 2023a; Chakraborty et al., 2024), 60% data is used as historical (training) data and the most recent 40%—consisting of 9,097 bugs—is used for evaluation.
Table 1. Ye et al. Dataset Description
Project # of Bug Reports # of Java Files LOC
Max Median Min Max Median Min
AspectJ 593 6,879 4,439 2,076 699,250 515,153 216,387
Birt 4,178 9,697 6,841 1,700 2,322,074 1,549,525 412,795
Eclipse 6,495 6,243 3,454 382 1,156,041 637,158 47,348
JDT 6,274 10,544 8,184 2,294 922,812 609,378 133,323
SWT 4,151 2,795 2,056 1,037 934,357 639,216 301,698
Tomcat 1,056 2,042 1,552 924 487,701 413,472 270,046
For RQ2 and RQ3, we use the GHRB (GitHub Recent Bugs) dataset, which is specifically designed for evaluating LLM-based approaches (Lee et al., 2024). It contains 131 recent bugs from 16 projects (e.g., Apache Dubbo, OpenAPI Generator), as shown in Table 2. The size of the dataset enables us to execute LLM-based approaches three times to account for their non-deterministic behavior. To further validate our findings, we also experiment with the latest 300 bugs (50 from each project) from the Ye et al. dataset. For RQ4, we use the subset of GHRB bug reports submitted after the knowledge cut-off date of GPT-4o-mini (October 1, 2023) (gpt, [n. d.]), leaving 49 unseen bug reports for evaluation.
Table 2. GHRB Dataset Description
# of Projects # of Bug Reports # of Java Files LOC
Max Median Min Max Median Min
16 131 12,025 1,977 93 1,874,139 299,697 14,759
4.5. Evaluation Metrics
Similar to other work (Zhou et al., 2012; Rahman and Roy, 2018), the following three metrics are used for evaluation:
1. Accuracy@k: It denotes the number of bugs for which at least one of the actual buggy files is ranked within the top-k positions (k = 1, 5, 10) of the ranked list. A higher accuracy@k value indicates a better performance of the bug localization technique. Similar to other studies (Qi et al., 2021; Liang et al., 2022), the value of k is set to 1, 5 and 10.
2. Mean Reciprocal Rank@k (MRR@k): This metric measures the average reciprocal rank of the first correctly identified buggy file within the top-k results across all bug reports. For each bug report, the reciprocal rank is defined as the inverse rank of the first correct buggy file that appears among the top-k results. MRR@k is calculated as follows:
(1) $MRR@k=\frac{1}{N}\sum_{i=1}^{N}\begin{cases}\frac{1}{\text{rank}_{i}}&\text{if }\text{rank}_{i}\leq k\\ 0&\text{otherwise}\end{cases}$
where, $N$ denotes the total number of bug reports. $\text{rank}_{i}$ refers to the rank of the first correct buggy file for the i-th bug report among the top-k results. If no actual buggy file appears within the top-k results for a bug report, the reciprocal rank is defined as 0. The higher the value of MRR@k, the better the technique. Following existing studies (Rahman and Roy, 2018; Kim and Lee, 2019), the value of k is set to 10. An MRR@10 value of 0.2 indicates that, on average, the first relevant file is ranked around position 5 within the top 10 results.
3. Mean Average Precision@k (MAP@k): This metric evaluates how well a bug localization technique ranks all actual buggy files within the top-k retrieved results for a set of bug reports. For each bug report, it calculates the Average Precision (AP) using the following formula, which considers both the presence and the ranking positions of all relevant buggy files:
(2) $AP_{i}=\frac{\sum_{j=1}^{M}P(j)\times\text{Rel}(j)}{\text{Number of relevant source files}}$
where, $M$ is the number of retrieved files in the top-k ranked list for the i-th bug report (i.e., MM=$k$). $Rel(j)$ denotes whether the file at position $j$ is relevant or not. $P(j)$ indicates the precision at rank $j$ and computed as follows:
(3) $P(j)=\frac{\text{Number of relevant files in top-}j\text{ positions}}{j}$
The Mean Average Precision is then computed by averaging the AP scores across all bug reports, as shown here:
(4) $MAP=\frac{\sum_{i=1}^{N}AP_{i}}{N}$
where, $N$ denotes the total number of bug reports. A higher MAP@k value indicates better ranking quality of the buggy files. Following prior studies (Rahman and Roy, 2018; Kim and Lee, 2019), the value of k is set to 10. For example, a MAP@10 value of 0.3 implies that, on average, 30% of the actual buggy files are retrieved with high precision within the top 10 results.
5. Result Analysis
In this section, the experimental results in accordance with the research questions are discussed.
5.1. RQ1: Comparison with Traditional and Deep Learning-based IRBL Techniques
GenLoc is evaluated against four traditional and deep-learning based baseline techniques (DreamLoc, BRTracer, BLUiR and BugLocator) across six widely-used benchmark projects (AspectJ, Birt, Eclipse, JDT, SWT and Tomcat). Tables 3 and 4 present the overall and project-wise results, respectively.
Table 3. Comparison of GenLoc and Non-LLM-based Bug Localization Techniques
Technique Accuracy@k (%) MAP@10 MRR@10
k=1 k=5 k=10
GenLoc 44.01 63.27 68.28 0.41 0.52
DreamLoc 26.33 52.32 63.07 0.30 0.37
BRTracer 26.84 51.21 60.42 0.29 0.37
BLUiR 26.49 49.04 59.82 0.28 0.36
BugLocator 24.19 46.20 56.65 0.26 0.33
Table 3 shows that GenLoc demonstrates strong overall performance across all evaluation metrics compared to prior bug localization techniques. On average, GenLoc achieves the highest Accuracy@1 (44.01%), substantially outperforming the second-best technique (BRTracer with 26.84%) by a relative margin of over 63%. This indicates that GenLoc can correctly identify the buggy file at the top position in a significant portion of bug reports. Similarly, GenLoc leads in Accuracy@5 (63.27%) and Accuracy@10 (68.28%), further affirming its ability to prioritize buggy files effectively within the top candidates. In terms of ranking quality, GenLoc achieves the highest scores in both MAP@10 (0.41) and MRR@10 (0.52), outperforming the next-best MAP@10 (DreamLoc with 0.30) and MRR@10 (BRTracer and DreamLoc with 0.37) by 36.67% and 40.54%, respectively, confirming its superiority in ranking relevant files near the top of the result list.
It is worth mentioning that GenLoc uses only the bug report as input, without incorporating any additional information such as historical bug-fix data or file change frequency. In contrast, techniques like DreamLoc, BRTracer and BugLocator leverage such additional data to improve localization accuracy. Despite considering less information, GenLoc consistently outperforms these baselines on all metrics.
Table 4. Project-wise Performance Comparison of Bug Localization Techniques
Project Technique Accuracy@k (%) MAP@10 MRR@10
k=1 k=5 k=10
Aspectj GenLoc 48.52 62.31 65.96 0.41 0.55
DreamLoc 14.77 39.66 52.74 0.18 0.26
BRTracer 25.32 55.27 64.56 0.29 0.38
BLUiR 24.05 42.19 58.22 0.23 0.33
BugLocator 18.57 39.24 51.90 0.20 0.28
Birt GenLoc 19.67 35.33 39.90 0.19 0.26
DreamLoc 19.33 39.38 49.13 0.21 0.28
BRTracer 14.00 28.67 37.52 0.14 0.21
BLUiR 13.88 29.80 37.88 0.14 0.20
BugLocator 12.39 26.93 34.95 0.13 0.17
Eclipse GenLoc 47.25 66.95 72.07 0.45 0.56
DreamLoc 34.72 62.36 72.06 0.37 0.46
BRTracer 29.37 54.66 63.74 0.32 0.40
BLUiR 25.52 48.61 57.97 0.28 0.35
BugLocator 28.37 52.12 61.86 0.31 0.38
JDT GenLoc 42.97 67.48 73.75 0.40 0.53
DreamLoc 29.21 54.64 64.48 0.31 0.40
BRTracer 27.62 57.37 64.05 0.28 0.38
BLUiR 30.61 55.08 64.81 0.30 0.41
BugLocator 25.35 51.02 62.38 0.27 0.36
SWT GenLoc 42.07 63.69 70.56 0.42 0.51
DreamLoc 28.92 58.13 69.64 0.34 0.41
BRTracer 19.94 42.35 54.94 0.25 0.30
BLUiR 25.06 51.75 63.98 0.29 0.36
BugLocator 19.94 39.64 51.08 0.23 0.29
Tomcat GenLoc 63.59 83.89 87.44 0.60 0.72
DreamLoc 31.04 59.72 70.38 0.36 0.43
BRTracer 44.79 68.96 77.73 0.46 0.55
BLUiR 39.81 66.82 76.07 0.42 0.51
BugLocator 40.52 68.25 77.73 0.44 0.52
Table 4 shows that GenLoc achieves the highest Accuracy@1 across all six projects. Notably, it demonstrates substantial improvements–ranging from 36% to 91%–over the best-performing baseline on projects such as AspectJ, Eclipse, JDT, SWT and Tomcat. At Accuracy@5, GenLoc continues to outperform all baselines in five out of six projects, most notably achieving 83.89% on Tomcat, 67.48% on JDT, and 66.95% on Eclipse. A similar trend is observed for Accuracy@10, where GenLoc continues to outperform other techniques.
Regarding ranking quality, GenLoc achieves the highest MAP@10 and MRR@10 scores on five of the six projects. For instance, on AspectJ, GenLoc achieves an MRR of 0.55, far exceeding that of the second-best technique (BRTracer with 0.38). Similar trends are observed in the MAP scores on projects such as Eclipse, JDT, SWT and Tomcat, where GenLoc consistently outperforms all baselines, further demonstrating its strength in producing effective ranked outputs.
Although GenLoc achieves the best Accuracy@1 (19.67%) on Birt project, its Accuracy@5 and Accuracy@10 are lower compared to DreamLoc. Additionally, GenLoc attains the second-highest MAP (0.19) and MRR (0.26), with DreamLoc slightly outperforming it on both metrics. To investigate this performance gap, the authors manually analyzed several Birt bug reports that GenLoc failed to localize and found a common pattern that many of these reports contained “steps to reproduce” information. While such descriptions are helpful for developers (Bettenburg et al., 2008; Soltani et al., 2020), they primarily focus on user actions (e.g., clicking buttons or selecting options), which rarely align directly with the source code, as shown in Fig. 3. This disconnection makes it difficult for GenLoc to extract meaningful semantic clues from the report and map them to relevant files.
Bug ID: 324546
Summary: external .js file has no effect in beforeOpen if the data set is used to generate a dynamic parameter selection
Description: Build Identifier: I define a function "replace_query" in an external .js file, and then include it in the report. I then use the function in the beforeOpen script. If the data set is not used to generate a dynamic parameter selection, then it works fine. Otherwise, when viewing the report, it will give this error: Error evaluating Javascript expression. Script engine error: ReferenceError: "replace_query" is not defined. Reproducible: Always Steps to Reproduce: 1. Define a function in an external .js file 2. Include the .js file in the report 3. Call the function in the beforeOpen script of a data set that is also used to generate a dynamic parameter selection.
Figure 3. Bug Report from Birt Project.
To systematically verify this observation, reproduction related keywords are extracted from all bug reports following prior work (Soltani et al., 2020). As shown in Table 5, 63.85% of Birt’s bug reports include such reproduction-related content, which is substantially higher than any other project (e.g., 3.80% in AspectJ, 2.62% in Eclipse, and 2.91% in JDT). This discrepancy contributes to the semantic mismatch between bug reports and source files, which hampers GenLoc’s ability to accurately rank relevant files. Furthermore, a chi-square test revealed a statistically significant relationship between the presence of reproduction information and the failure to localize bugs in the Birt project across all three GenLoc trials (p-value = 0). In contrast, DreamLoc considers additional features such as bug-fixing history of a file, cyclomatic complexity, apart from similarity matching. These supplementary features helped DreamLoc outperform GenLoc on Birt. However, when DreamLoc is executed without these additional features, its performance also drops significantly, as shown in Table LABEL:dreamloc.
Table 5. Distribution of bug reports with reproduction-related keywords
Project Total Bugs Reproduction-related Keywords Percentage (%)
AspectJ 237 9 3.80
Birt 1671 1067 63.85
Eclipse 2598 68 2.62
JDT 2509 73 2.91
SWT 1660 88 5.30
Tomcat 422 22 5.21
Table 6. Comparison of GenLoc and DreamLoc (without additional features) on the Birt Project
Technique Accuracy@k (%) MAP@10 MRR@10
k=1 k=5 k=10
GenLoc 19.67 35.33 39.90 0.19 0.26
DreamLoc (w/o feats) 4.73 11.97 16.16 0.08 0.05
Figure 4. Overlap Analysis between GenLoc and Non-LLM based IRBL techniques.
Apart from Accuracy@k, MRR@10 and MAP@10, the number of unique bugs localized by each approach are examined. In this analysis, a bug is considered successfully localized if the correct buggy file appears within the top 10 ranked results (i.e., Accuracy@10) (Kochhar et al., 2016). Figure 4 presents the unique bugs localized by each technique under this criterion. Since GenLoc is a non-deterministic approach, its output varies across different executions due to the inherent randomness in language model sampling. To fairly evaluate GenLoc’s overall capability and compare it with existing deterministic baselines, the union of all bugs localized across three trials is considered. This union represents the total set of unique bugs that GenLoc is able to localize in at least one run, capturing the full potential of the approach rather than the outcome of a single arbitrary execution.
Among all approaches, GenLoc can localize the highest number of unique bugs (601 bugs). This indicates that GenLoc not only localizes more bugs overall but also captures a unique subset of bugs missed by all other methods, highlighting its potential to complement existing methods. This is likely due to its ability to reason over natural language and semantic context, which prior techniques lack.
5.2. RQ2: Comparison with LLM-based IRBL Techniques
Table 7 reports the comparison of GenLoc against three LLM-based techniques—BRaIn, LLM-BL and Agentless—using the GHRB dataset. The results show that GenLoc consistently outperforms all baselines across every evaluation metric. It achieves the highest Accuracy@1 of 63.36%, substantially higher than other techniques, and maintains this advantage at k=5 and k=10. For ranking-based measures, GenLoc achieves a MAP@10 of 0.59 and MRR@10 of 0.69, about 20% higher than LLM-BL, while outperforming Agentless by over 75% and BRaIn by more than 100%. In addition, GenLoc localizes 5 unique bugs (Fig. 5), demonstrating that it can complement other LLM-based techniques. We observed similar results on the subset of Ye et al. dataset, with detailed results provided in the replication package.
Table 7. Performance of LLM-based IRBL Techniques
Technique Accuracy@k (%) MAP@10 MRR@10
k=1 k=5 k=10
GenLoc 63.36 76.85 79.65 0.59 0.69
BRaIn 22.90 51.15 61.07 0.28 0.34
LLM-BL 50.38 63.11 67.18 0.47 0.56
Agentless 28.50 52.93 65.40 0.33 0.39
Figure 5. Overlap Analysis between GenLoc and Recent LLM-based Approaches.
In Fig. 6, representative lines from a bug report [^1] https://github.com/apache/rocketmq/issues/7026 are shown where GenLoc could localize the correct bug file within top 2 rank across all trials, whereas no other technique could localize it even within top 5. The report noted that the proxy returned a null address when the master broker was down instead of failing over to a slave. GenLoc first applied embedding-based retrieval to narrow the search space, where ClusterTopicRouteService.java aligned with the bug context. It then analyzed the method signatures of this file to confirm its relevance and finally inspected method bodies to uncover the missing failover logic. In contrast, neither BRaIn nor LLM-BL could shortlist the correct file, and Agentless, despite exposing the entire repository structure to the LLM, failed because the correct file was lost in the overwhelming directory and file information. By contrast, GenLoc’s two-stage process placed the correct file within the Top-50 through embeddings and then promoted it to the Top-2 through LLM based reasoning.
Bug ID: 7027
Summary: [Bug] The proxy in the cluster mode returns null adress when master is down
Description:Describe the Bug When I connect to the clustered proxy to consume messages using the remoting protocol, the master broker happened to hang up. At this time, the address returned by the proxy was empty, which caused the request to be sent to the NameServer. Steps to Reproduce 1. Deploy tow group of brokers, such as: 1. master: broker-a, slave: broker-a-s 2. master: broker-b, slave: broker-b-s 2. Deploy a proxy 3. Create some topics in broker-a for testing, such as message produing and consuming, that is not important. 4. **Kill master broker-a**. # What Did You Expect to See? Proxy should return slave address, for consumer can continue consume message from slave. What Did You See Instead? Proxy send request to NameServe. # Additional Context _No response_
Figure 6. Bug Report from Apache RocketMQ Project.
5.3. RQ3: Ablation Study
To better understand the contribution of different components in GenLoc’s bug localization pipeline, we conduct an ablation study using the following four configurations:
  1. (1)
    GenLoc: This is the complete version of the proposed approach, where the LLM is provided with five functions, including get_candidate_filenames() to utilize embedding-based retrieval to identify potentially buggy files.
  2. (2)
    Embedding-Only: A baseline that evaluates the standalone performance of the embedding-based retrieval system, where source files are ranked solely based on their semantic similarity to the bug report, without any LLM reasoning or iterative refinement.
  3. (3)
    GenLoc-NoEmbed: A variant where the LLM operates without access to the embedding-based recommended files, i.e., the get_candidate_filenames() function is removed. The model must infer filenames directly from the bug report and continue reasoning with the remaining four functions.
  4. (4)
    GenLoc-Naive: A minimal setup, where the LLM is limited to using only two functions—search_file() and search_method()—and must rely solely on basic keyword-based exploration without access to embedding guidance or deeper code-level analysis.
Table 8. Result of Ablation Study
Technique Accuracy@k (%) MAP@10 MRR@10
k=1 k=5 k=10
GenLoc 63.36 76.85 79.65 0.59 0.69
Embedding-Only 20.36 58.27 72.52 0.30 0.37
GenLoc-NoEmbed 46.05 56.24 57.76 0.42 0.50
GenLoc-Naive 46.31 57.25 58.78 0.43 0.50
Table 8 presents the average results on the GHRB dataset across all four configurations. We observed similar results on the subset of Ye et al. dataset, which can be found in the replication package. Results show that GenLoc achieves the best performance on all metrics, demonstrating the strength of combining embedding-based retrieval with iterative LLM analysis. The Embedding-Only variant obtains only 20.36% at Accuracy@1, indicating that retrieval alone is insufficient without deeper analysis. Similarly, the GenLoc-NoEmbed achieves only 46.05% Accuracy@1, showing that the absence of embedding guidance limits the LLM’s ability to identify relevant files. These results highlight that embedding-based retrieval and LLM-based reasoning are complementary, and removing either degrades localization performance.
The performance improvement of GenLoc over GenLoc-NoEmbed and GenLoc-Naive indicates that GenLoc goes beyond superficial cues (e.g., filenames or keywords) and performs deeper reasoning. A case [^2] https://github.com/OpenAPITools/openapi-generator/issues/19039 is shown in Fig. 7, where GenLoc-NoEmbed and GenLoc-Naive failed to localize the correct file, while GenLoc succeeded. The bug report described an error in the TypeScript Fetch generator of OpenAPI, where the modelNamePrefix option was being incorrectly applied twice in generated import statements. Misled by surface-level clues, GenLoc-NoEmbed and GenLoc-Naive attempted to locate a nonexistent file named ModelNamePrefix and also searched for the generate() method, which was unrelated to the bug and provided no useful signal. In contrast, GenLoc invoked get_candidate_filenames() to narrow down to the TypeScript-specific implementation and by analyzing method signatures and bodies in TypeScriptFetchClientCodegen.java (e.g., toModelFilename(), parseImports()), it successfully localized the buggy file.
Bug ID: 20109
Summary: [BUG] [typescript-fetch] ModelNamePrefix is added twice to import statements thereby breaking the build
Description: When generating the model files the prefix is properly added once in front of the generated files. However, in the import statements the prefix is added _twice_ resulting in erroneous statements breaking the build. When e.g. ‘SomePrefix‘ is set for ‘modelNamePrefix‘, then the following model files are generated for ‘example-for-file-naming-option.yaml‘: * ‘SomePrefixPetCategory.ts‘ * ‘SomePrefixPet.ts‘ However, _within_ ‘SomePrefixPet.ts‘ it looks like this: “‘typescript import type SomePrefixPetCategory from ’./SomePrefixSomePrefixPetCategory’; import SomePrefixPetCategoryFromJSON, SomePrefixPetCategoryFromJSONTyped, SomePrefixPetCategoryToJSON, from ’./SomePrefixSomePrefixPetCategory’; “‘ I.e. ‘SomePrefix‘ is added twice.
Figure 7. Bug Report from OpenAPI Generator Project.
Figure 8. Overlap analysis between GenLoc and Its Ablated Variants.
Although GenLoc-NoEmbed and GenLoc-Naive obtain similar results, the overlap analysis (Fig. 8) shows that they captures distinct subsets of bugs. This demonstrates the importance of combining all four functions—search_file(), search_method(), get_method_signatures_of_a_file() and get_method_body()—to achieve broader coverage. Their integration enables GenLoc to leverage both keyword-level and structure-aware reasoning.
Finally, Table 9 presents the consistency of results, measured as consensus across trials for each variant (i.e., GenLoc, GenLoc-NoEmbed and GenLoc-Naive). Here, consensus indicates the proportion of bugs that are localized in at least two or all three trials, reflecting the stability of each approach. GenLoc shows the highest stability, with 93.69% of bugs localized in at least two runs and 88.29% in all three runs. In contrast, GenLoc-NoEmbed and GenLoc-Naive attain lower values, reflecting greater variability across trials. It shows that the integration of all components not only improves accuracy but also enhances the consistency of GenLoc’s results.
Table 9. Consistency Across Trials
Consensus GenLoc (%) GenLoc-NoEmbed (%) GenLoc-Naive (%)
$\geq$ 2 trials 93.69 85.23 86.36
$\geq$ 3 trials 88.29 72.73 76.14
5.4. RQ4: Performance on Unseen Bugs
Table 10 presents the performance of GenLoc on both seen and unseen bug reports from the GHRB dataset. The results indicate that GenLoc maintains consistent performance across the two subsets. For instance, the accuracy@1 remains stable at approximately 63% for both seen and unseen bugs. To further validate this observation, we performed a Mann–Whitney U test, a non-parametric method suitable for comparing independent groups (McKnight and Najab, 2010). The results reveal no statistically significant difference in Reciprocal Rank (RR) (p-value > 0.7) and Average Precision (AP) (p-value > 0.3) between the two groups. These findings suggest that GenLoc does not rely on memorization or training data leakage and can effectively localize bugs even in reports that fall outside the temporal scope of the LLM’s training data.
Table 10. Performance on Seen and Unseen Bugs
Category Accuracy@k (%) MAP@10 MRR@10
k=1 k=5 k=10
Seen 63.41 78.05 80.90 0.60 0.70
Unseen 63.27 74.83 77.55 0.56 0.68
6. Discussion
The evaluation demonstrates that GenLoc consistently outperforms traditional IR-based, deep learning–based and recent LLM-based bug localization techniques. Its strength lies in integrating semantic retrieval with LLM-driven iterative reasoning, which ensures that relevant files are surfaced and analyzed without overwhelming the model. This design highlights the importance of context quality in LLM-driven bug localization—too little context leads to irrelevant analysis, while too much hinders reasoning. By striking this balance through embedding-based retrieval and targeted exploration, GenLoc achieves reliable and consistent performance improvements across diverse projects.
Apart from higher accuracy, GenLoc is also cost-effective. On the Ye et al. dataset, the average cost to localize a bug is \$0.010, which includes \$0.003 for embedding-based retrieval (Step 1) and \$0.007 for LLM-based reasoning (Step 2). On the more recent GHRB dataset, the cost is \$0.025 per bug, with \$0.019 for embeddings and \$0.006 for LLM reasoning. The lower per-bug cost on the Ye et al. projects, despite their larger size in terms of files and lines of code, arises from their multi-version structure: once the initial embedding is completed, only changed or newly added files need to be reprocessed in later versions. By contrast, GHRB includes more projects but far fewer bug reports per project, so a larger fraction of projects must go through the full initial embedding stage without enough subsequent bugs to offset this overhead. Nevertheless, the per-bug cost of \$0.025 remains reasonable, demonstrating that GenLoc is affordable and practical for both long-lived projects and newer projects with fewer versions. Moreover, unlike BugLocator, BRTracer and DreamLoc, GenLoc relies solely on the bug report, thereby making it broadly applicable even in projects without historical bug-fix data.
In terms of efficiency, GenLoc can produce results within practical time limits. The average processing time per bug is 48.35 seconds on the Ye et al. dataset and 37.76 seconds on GHRB. These times are consistent with findings from a prior study (Kochhar et al., 2016) which reported that 90% of developers expect a bug localization technique to return results within one minute. This alignment suggests that GenLoc can operate within developer expected timeframe while also maintaining high accuracy and low cost, highlighting its potential suitability for real-world debugging scenarios.
Despite its success, GenLoc has several limitations. It may fail to localize bugs when the bug report lacks semantic or behavioral clues—particularly when the report is written from a user perspective and focuses mainly on reproduction steps without mentioning relevant program behavior or entities. One promising direction to address this challenge is to integrate an LLM-driven query reformulation step that rewrites or enriches the original bug report to better align with the source code and improve the quality of retrieved candidates.
Another limitation is that GenLoc may not identify all files associated with a given bug within its top-10 ranked list, which is the current output size of the system. However, a prior work from Facebook (Pradel et al., 2020) suggests that predicting even a single relevant file is useful in practice since it helps route the issue to the appropriate developer or team and can serve as a starting point for discovering additional affected files through manual inspection or complementary techniques. Although this does not replace the need for comprehensive localization, it highlights the potential value of partial predictions in real-world debugging workflows.
7. Threats to Validity
This section presents potential ways in which the validity of the study may be compromised.
Internal Validity: One of the most significant threats to internal validity is the possibility of data leakage, such as bug-fixing commits being present in the LLM’s training data. However, results of RQ4 show that GenLoc maintains consistent performance on both seen and unseen bug reports in the GHRB dataset, indicating that GenLoc effectively localizes bugs even for reports outside the temporal scope of the LLM’s training data. Another potential threat arises from the inherent randomness of LLMs, which may yield different results across different runs. To mitigate this, all experiments were repeated three times, and a replication package is released to support reproducibility and enable further validation by the research community. Additionally, the choice of LLM may impact the results, as different models can vary significantly in reasoning capabilities, code understanding and sensitivity to prompt structure. Hence, future evaluations should assess the robustness of GenLoc across multiple LLM variants.
Construct Validity: Construct validity concerns whether the evaluation metrics accurately measure the effectiveness of bug localization techniques (Widyasari et al., 2022). This study employs Accuracy@k, MRR@10 and MAP@10, which are widely used in prior bug localization research and considered standard for assessing ranking quality (Wong and Tse, 2023; Rahman and Roy, 2018).
External Validity: GenLoc is evaluated on two complementary datasets: a large benchmark of over 9,000 bug reports from six open-source Java projects that has been widely used in prior bug localization research (Chakraborty et al., 2024), and the more recent GHRB dataset of 131 bugs from 16 projects, designed specifically for evaluating LLM-based approaches (Lee et al., 2024). While these datasets provide strong evidence of GenLoc’s effectiveness, the findings may still not generalize to proprietary industrial systems or projects written in other programming languages. Additional studies are needed to evaluate the effectiveness of GenLoc in broader and more heterogeneous settings.
8. Related Work
This section discusses papers from two related domains.
8.1. Information Retrieval-based Bug Localization
Existing IRBL approaches can be broadly categorized into three groups based on their underlying methodology. The first category (Lukins et al., 2008; Nguyen et al., 2011; Zhou et al., 2012; Sisman and Kak, 2012; Kim et al., 2013; Saha et al., 2013; Wong et al., 2014) uses traditional IR models such as Latent Dirichlet Allocation (LDA) and VSM for bug localization. While Lukins et al. were the first to apply LDA for bug localization (Lukins et al., 2008), BugScout extended this idea by adding bug history and file size (Nguyen et al., 2011). BugLocator introduced a revised VSM incorporating file size and similar past bugs (Zhou et al., 2012). BLUiR further improved retrieval by using structured code elements such as class and method names (Saha et al., 2013), while AmalGAM combined the strengths of BugLocator and BLUiR with version history (Wang and Lo, 2014). Sisman and Kak introduced bug and file modification history (Sisman and Kak, 2012), and Kim et al. improved precision by filtering uninformative bug reports (Kim et al., 2013). BRTracer included stack traces and code chunking (Wong et al., 2014), while later efforts such as Locus (Wen et al., 2016) and BLIA (Youm et al., 2017) incorporated fine-grained change information.
The second category (Ye et al., 2014; Fejzer et al., 2021; Lam et al., 2015, 2017; Xiao et al., 2019; Cao et al., 2020; Qi et al., 2021; Wang et al., 2020; Iqbal et al., 2020; Liang et al., 2022; Ciborowska and Damevski, 2022; Du and Yu, 2023; Chakraborty et al., 2024; Luo et al., 2022; Anh and Luyen, 2021; Zakari et al., 2020) consists of learning-based models that apply machine learning and deep learning to capture richer semantic relationships between bug reports and source code. (Ye et al., 2014; Fejzer et al., 2021) leveraged learning-to-rank model for bug localization. Early deep learning methods such as HyLoc (Lam et al., 2015), DNNLOC (Lam et al., 2017), and DeepLoc (Xiao et al., 2019) used DNNs to find semantic similarity between bug report and source code. CNN-based models like BugPecker (Cao et al., 2020), DreamLoc (Qi et al., 2021), MD-CNN (Wang et al., 2020), and TRANP-CNN (Iqbal et al., 2020) enhanced structural understanding using ASTs, CFGs, and transfer learning. Later approaches, including FLIM (Liang et al., 2022), FBL-BERT (Ciborowska and Damevski, 2022), HMCBL (Du and Yu, 2023) and RLocator (Chakraborty et al., 2024) leveraged pre-trained language models as feature extractors to improve semantic matching and scalability. Other models like CoLoc (Luo et al., 2022), ImbalancedBugLoc (Anh and Luyen, 2021), and DependLoc (Zakari et al., 2020) addressed specific challenges such as class imbalance, cross-project localization, and dependency-aware analysis.
The third category consists of reasoning-enhanced approaches (Li et al., 2025b; Samir and Rahman, 2025; Li et al., 2025a) where language models are used as an active reasoning component. KEPT enhances the UniXcoder language model by incorporating project-specific knowledge graphs, and fine-tuning it using contrastive learning (Li et al., 2025b). BRaIn combines lexical retrieval with LLM-based method analysis and iterative refinement (Samir and Rahman, 2025), while LLM-BL employs LLM-driven query expansion and hybrid retrieval to enhance the ranking of buggy files and lines (Li et al., 2025a). GenLoc differs from prior LLM-based approaches by combining semantic retrieval with iterative code-exploration functions, allowing the LLM to selectively inspect and reason over the code base rather than relying solely on retrieval or reranking.
8.2. LLMs in Software Debugging and Automated Issue Resolution
Recent studies have investigated the usage of LLMs in software debugging and automated issue resolution. A notable line of work focuses on spectrum-based bug localization, where approaches such as AutoFL (Kang et al., 2024), LLM4FL (Rafi et al., 2024), CosFL (Qin et al., 2025a) and SoapFL (Qin et al., 2025b) leverage failing test cases to identify faulty code segments. Building on this, FuseFL (Widyasari et al., 2024) enhances interpretability by generating step-by-step reasoning to justify faulty lines, while FlexFL (Xu et al., 2025) further improves accuracy by combining bug reports with execution traces.
In parallel, other research efforts aim to enhance bug reports using LLMs. For instance, ChatBR (Bo et al., 2024) generates missing information such as observed and expected behavior, while (Kang et al., 2023) and (Plein and Bissyandé, 2023) investigate how LLMs can generate test cases directly from bug reports. However, none of these studies analyzed the impact of LLM in the context of IRBL, when no other information (e.g. test cases) is available.
Another emerging direction involves the use of LLMs for automated issue resolution on platforms like GitHub. Techniques such as MAGIS (Tao et al., 2024), AutoCodeRover (Zhang et al., 2024), SpecRover (Ruan et al., 2025), LingmaAgent (Ma et al., 2025) and Agentless (Xia et al., 2025) aim to address software issues including bugs. However, these methods primarily focus on the repair phase, often overlooking the bug localization step, which remains a crucial and challenging part of the debugging process (Chang et al., 2025). In contrast, this paper focuses on the bug localization phase, investigating how LLMs can be leveraged to enhance IRBL.
9. Conclusion and Future Work
This paper presents GenLoc, a novel information retrieval-based bug localization technique that integrates semantic retrieval with step-by-step reasoning capabilities of LLMs. An evaluation on over 9,000 real-world bugs from six diverse Java projects and the GHRB dataset of 131 recent bugs across 16 projects shows that GenLoc consistently outperforms traditional IR-based methods, deep learning models and recent LLM-based techniques in terms of Accuracy@k, MRR@10 and MAP@10. Moreover, GenLoc can localize the highest number of unique bugs, demonstrating its potential to complement existing approaches.
In the future, GenLoc could be extended to support finer-grained bug localization at the method or statement level to provide more precise debugging assistance. Another promising direction is to integrate GenLoc with automated program repair pipelines to evaluate its effectiveness in guiding correct patch generation.